### v1

# Fast and Accurate Decoder for the XZZX code Using Simulated Annealing

Abstract
The XZZX code is a variant of the surface code designed to address biased noise in realistic quantum devices. For the XZZX code, we propose a decoder based on simulated annealing (SA). Our SA decoder can be readily and efficiently parallelized, by virtue of its simple MCMC-based algorithm. To prepare an initial configuration of SA, we propose to employ recovery chains obtained by a decoder which utilizes a kind of greedy matching graph algorithm. Although $Z$-biased noise is commonly assumed in real quantum devices, we focus on $Y$-biased noise, for which the minimum-weight perfect matching (MWPM) algorithm fails to decode accurately. Our numerical simulation for the code capacity noise model, where only data qubits suffer errors, confirmed that our SA decoder is more accurate than the MWPM decoder. Furthermore, our SA decoder attained the accuracy equivalent to that of the optimal decoder formulated by integer programming, called CPLEX decoder. In our greedy matching decoder, we randomly determine order of matching pairs of incorrect syndromes that have the same distance. This randomness brings about a variety of initial configurations of SA, which leads to faster convergence of our SA decoder. By comparing decoding times of our SA decoder, the CPLEX decoder, and matrix product state (MPS) decoder, all of which can handle $Y$-biased noise appropriately, we confirmed that our SA decoder is fastest if parallelized. This result implies a potential for combining of our greedy matching and SA decoder for practical use in quantum computing.
XZZX code, simulated annealing, biased noise, code capacity noise, maximum likelihood decoder, Metropolis algorithm
I Introduction
The ultimate goal of quantum computing is to realize fault-tolerant quantum computation (FTQC) with robust error correction mechanisms. For superconducting devices, decoherence time (qubit lifetime) is typically on the order of hundreds of microseconds. For this reason, the decoding process must be completed within this time frame. Practical decoders require a simple algorithm to be readily implemented in hardware.
The surface code is one of the most prospective topological error correction codes [1, 2]. It is experimentally feasible, because it requires measuring only nearest-neighbor data qubits [3, 4, 5, 6, 7].
In physical systems such as superconducting qubits, optics and ion traps, realistic noise is biased, that is, $Z$ (phase-flip) errors occur more frequently than $X$ (bit-flip) errors. For such a biased noise, replacing the stabilizer generators of the surface code leads to alternative codes, such as the tailored surface code (which has only $Y$-type stabilizers) [8] and the XZZX code [9]. The XZZX code exhibits higher thresholds for $X$-,$Y$- or $Z$-biased noises, compared with the ordinary surface code, which has stabilizer generators consisting of either only $X$ or $Z$ operators, and is known as the CSS (Calderbank–Shor–Steane) surface code [10, 11]. For the XZZX code, there appeared proposals on its rotated version [12, 13], its cluster state [14] and its physical realization in optical and superconducting devices [15] and neutral atom devices [16].
The minimum-weight perfect matching (MWPM) decoder [1] can be executed in polynomial time of the number of incorrect syndromes and, hence, is well used for the surface code. However, it deals with $X$ and $Z$ errors independently, and is thus less accurate under noises that include $Y$ errors.
On the other hand, MCMC (Markov chain Monte Carlo) decoder and matrix product state (MPS) decoder [17], known as MLD (maximum likelihood decoder), possess high decoding accuracy by taking into account the probability of $Y$ errors. For the $Z$-biased and $Y$-biased noises, the MPS decoder has been researched [18]. Unfortunately, the MPS decoder has a high computational complexity of order $\mathcal{O}(d^{2})$ for the code distance $d$. In contrast, the MCMC decoders have a possibility for fast decoding, since the MCMC algorithms use heuristic approach based on pseudo-random numbers.
One of the earliest studies [19] on MCMC decoders for the CSS surface code employed parallel tempering, while another study [20] utilized the Metropolis method fixed at the Nishimori temperature. For initial configuration of MCMC, they employ error configurations generated by acting stabilizers and logical operators randomly to actual error configuration [19, 12] and recovery chain obtained by the MWPM decoder [20]. Thus, the aim of these previous works is to establish an accurate threshold for the code.
To efficiently minimize energy in MCMC methods, simulated annealing (SA) [21] is often used in statistical physics. By formulating the decoding problem as a random-bond Ising model (RBIM), SA has been applied to decoders for the CSS surface code [22] and the color code [23]. Another formulation of SA decoder for the CSS surface code has hamiltonian with penalty terms [24], and is demonstrated in hardware-implemented annealing machine [25, 26, 22].
The XZZX code has been studied by leveraging the MWPM decoder, the MPS decoder [9] and MCMC decoder [12]. They investigated the decoding accuracy for $Z$-biased noise, which is typical in real quantum devices. Assuming knowing the probabilities of $X,Y,Z$ errors, they incorporate this prior information into weight coefficients of the hamiltonian so as to improve decoding accuracy.
The research [12] assumes the code capacity noise model, in which only data qubits suffer errors independently and identically. To explore configurations efficiently, they carried out the MCMC method at lower inverse temperature than the Nishimori temperature, keeping the weight coefficients invariant in the hamiltonian. Their numerical experiment suggests that calculating a dominant term (composed of degeneracy and energy) of the partition function is sufficient to determine the correct logical operator class. Thus, their algorithm may be useful to collect this term to utilize it as train data in neural network decoder.
In the present article, we propose a simulated annealing (SA) decoder for the XZZX code. The aim of our research is not to determine an accurate threshold, but rather to establish a decoding algorithm that is moderately accurate and yet sufficiently fast for practical use on a quantum computer. We assume the code capacity noise. Specifically, our interest is the case that each qubit experiences $Y$-biased noise, which cannot be decoded appropriately by the MWPM decoder. As initial configuration for SA, we propose to employ a recovery chain obtained from a decoder which utilizes a greedy matching algorithm. We implement the whole algorithms in the C++ language from scratch, in order to attain decoding time as short as possible.
The rest of the present paper is organized as follows. In Sec. II, we review the XZZX code. For the XZZX code, we describe principle of error detection and correction in Sec. III and decoding algorithms, including our SA decoders in Sec. IV. We show numerical result and argue decoding accuracy and time in Sec. V and discuss the benefits of our SA decoder in Sec. VI. In Sec. VII we conclude.
II The XZZX code
In this section, we briefly review the XZZX surface code [9]. We consider a rotated square lattice with open boundaries [27]. In the lattice, data qubits reside on the vertices.
(a) Refer to caption
(b) Refer to caption(c) Refer to caption
Figure 1: The XZZX code for the code distance $d=5$. (a) Stabilizer generators $G_{f}$, showing four-body in the bulk and three-body operators at the boundaries. (b) Logical operator $L_{Z}$ (c) Logical operator $L_{X}.$
We illustrate stabilizer generators for the XZZX surface code in Fig. 1(1). The stabilizer generator acting around a face $f$ is defined as
$G_{f}\coloneqq X_{\mathrm{left}(f)}\,Z_{\mathrm{up}(f)}\,Z_{\mathrm{down}(f)}\,X_{\mathrm{right}(f)}$ (1)
whose form is the origin of the name ‘XZZX code’. This stabilizer generator is a four-body operator in bulk and a three-body operator at the boundaries. The XZZX code has only one type of stabilizer generators, unlike the CSS surface code which has two types of generators. Since each of the stabilizer generator includes both $X$ and $Z$ operators as shown in Eq. (1), the XZZX code is not CSS-type. The number of stabilizer generators is equal to the number of faces, i.e., $N_{\mathrm{stabilizer}}=2d(d-1)=\mathcal{O}(d^{2})$.
At the center of every face $f$, an ancilla qubit is located to detect errors. For each ancilla qubit, we measure the stabilizer generator $G_{f}$.
We define logical operators $L_{Z}$ ($L_{X}$) as a chain connecting the left and right (upper and lower) boundaries, respectively. Fig. 1(1),(1) depict examples of the logical operators $L_{Z}$ and $L_{X}$. Furthermore, we define logical operator $L_{Y}\coloneqq\mathrm{i}\,L_{Z}\,L_{X}$. As long as the opposite boundaries are connected, we regard these chains as logical operators, namely, allow the freedom of a stabilizer operator in the stabilizer group $\mathcal{G}\coloneqq\langle G_{f}\,|\,f:\text{face}\rangle$. More precisely, for $P\in\{I,X,Z,Y\}$, the logical operator is defined as equivalence class $L_{P}]$ by imposing invariance under the action of the stabilizer group $\mathcal{G}$ in the sense that $G\,L_{P}]=[L_{P}]$ for any $G\in\mathcal{G}$.
The code distance $d$ is defined as the length of the shortest representation of the logical operators $L_{Z}$ and $L_{X}$. For the surface codes including the XZZX code, it is the linear length of the lattice. The number of data qubits is $N_{\mathrm{q}}=d^{2}+(d-1)^{2}$.
The XZZX code is advantageous under biased noises. That is, the XZZX code has threshold $p_{\mathrm{th}}=0.5$ for pure $X$, $Y$, $Z$ noises, because it is equivalent to classical repetition codes; On the other hand, the CSS surface code has $p_{\mathrm{th}}=0.109$ for the pure $X,Z$ noises [1] and $p_{\mathrm{th}}=0.5$ for the pure $Y$ noise [8].
III Principle of Error detection and correction
Throughout the present paper, error configuration means the information that each data qubit suffers which $I$ (no error), $X$, $Z$ or $Y$ error. Fig. 2 illustrates an example of the error configuration. In Fig. 2, X,Y,ZX,\,Y,\,Z errors occur on specific data qubits.
To describe error detection, we introduce two decoding graphs $(\mathcal{V}_{\mathrm{white}},\mathcal{E}_{\mathrm{white}})$ and $(\mathcal{V}_{\mathrm{gray}},\mathcal{E}_{\mathrm{gray}})$. We write $(\mathcal{V}_{\bullet},\mathcal{E}_{\bullet})$ with $\bullet\in\{\mathrm{white},\mathrm{gray}\}$ and define it as follows:
The set of vertices $\mathcal{V}_{\bullet}$
The vertices are the centers of faces with the color $\bullet$ and are the ancillas for the stabilizer measurement.
The set of edges $\mathcal{E}_{\bullet}$
The edges which connect nearby vertices in $\mathcal{V}_{\bullet}$. The edge centers represent errors which suffer data qubits.
The set of edges which suffer errors is called error chain. There are two types of error chains $C_{\mathrm{white}}\subset\mathcal{E}_{\mathrm{white}}$ and $C_{\mathrm{gray}}\subset\mathcal{E}_{\mathrm{gray}}$, which connect centers of white and gray face, respectively, drawn with the pink and purple lines in Fig. 2. For both the error chains $C_{\mathrm{white}}$ and $C_{\mathrm{gray}}$, the horizontal (vertical) edges represent $Z$ ($X$) errors, respectively. Solely on the intersection points of the horizontal and vertical edges in the different types of error chains, $Y$ errors appear.
Figure 2: An example of the error configuration. The circles indicate qubits which suffer errors with error type $X$,$Y$,$Z$. (The qubits with no errors are not drawn.) The error configuration is expressed as two types of error chains $C_{\mathrm{white}}$ and $C_{\mathrm{gray}}$, connecting the white faces and gray faces and drawn with the pink and purple lines, respectively. The endpoints of these error chains are captured by the incorrect syndromes, drawn with the green squares. In both the error chains $C_{\mathrm{white}}$ and $C_{\mathrm{gray}}$, the $Z$ ($X$) errors reside on horizontal (vertical) edges, respectively. The $Y$ error appears solely on the intersection point of both the error chains $C_{\mathrm{white}}$ and $C_{\mathrm{gray}}$.
For each face $f$, we measure the stabilizer generator $G_{f}$ and thereby obtain a measurement outcome $s_{f}\in\{0,1\}$. For an error chain $C_{\bullet}$, it holds the eigenvalue equation
[…]
In Fig. 2, incorrect syndromes are drawn with green squares. The incorrect syndromes capture the endpoints of $C_{\mathrm{white}}$ and $C_{\mathrm{gray}}$. In the endpoint of horizontal (vertical) edges, $Z$ ($X$) errors anti-commute with $X$ ($Z$) operators acting on the left and right (up and down) qubits in the stabilizer generator $G_{f}$ and hence can be detected.
For an error chain $C_{\bullet}$, the set of face centers whose ancillas possess incorrect syndromes is called syndrome and denoted by
$S_{\bullet}\coloneqq\left\{(\text{center of face $f$})\in\mathcal{V}_{\bullet}\,|\,s_{f}=1\right\}.$ (3)
The syndrome consistency can then be expressed as
$\partial{C}_{\mathrm{white}}=S_{\mathrm{white}},\qquad\partial{C}_{\mathrm{gray}}=S_{\mathrm{gray}}.$ (4)
The two error chains $C_{\mathrm{white}}$ and $C_{\mathrm{gray}}$ contain the same information as the error configuration CC [^1] In the present article, to distinguish an error configuration $C$, we denote an error chain by $C_{\bullet}$ with $\bullet\in\{\mathrm{white},\mathrm{gray}\}$. . We express the correspondence as $C=(C_{\mathrm{white}},C_{\mathrm{gray}})$ and then define the tuple $\partial{C}\coloneqq(\partial{C_{\mathrm{white}}},\partial{C_{\mathrm{gray}}})$. Similarly, we define the tuple $S\coloneqq(S_{\mathrm{white}},S_{\mathrm{gray}})$. Using this abbreviated notation, the syndrome consistency of Eq. (4) reads $\partial{C}=S$.
The actual error configuration $C$ is unknown; The only information available is the syndrome $S$, from which decoders infer a recovery configuration $C$ such that $\partial{C}=S$. Decoders can be classified into two types: those that identify the maximum likelihood error (MLE) configuration and those that identify the maximum-likelihood logical operator class. The latter is referred to as the maximum likelihood decoder (MLD). The examples of the MLE decoder are the MWPM decoder (Sec. IV.2) and our greedy matching decoder (Sec. IV.3). The examples of MLD are our SA decoders (Sec. IV.1, Sec. IV.4), decoder formulated as integer programming (Sec. A) and the MPS decoder [17]. In general, the second type is more accurate than the first type, because it can take into account the degeneracy of error configurations with the same energy.
IV Decoding algorithms
We elaborate on our SA decoder in Sec. IV.1. Subsequently, we describe the MWPM decoder in Sec. IV.2. As a simplified version, in Sec. IV.3, we propose a greedy matching decoder, exploited to construct an initial configuration of the SA decoder.
IV.1 Simulated annealing (SA) decoder
We assume the code capacity noise where only the data qubits suffer errors. Specifically, each data qubit suffers an error of $X$-, $Y$-, or $Z$-type, independently and identically obeying the probabilities $p_{x}$, $p_{y}$, $p_{z}$, respectively. Any type of error occurs with the probability $p\coloneqq p_{x}+p_{y}+p_{z}$, called physical error probability. This noise for each data qubit $\rho$ is expressed as quantum channel
$\mathcal{E}(\rho)=(1-p)\,\rho\,+\,p_{x}X\rho X+p_{y}Y\rho Y+p_{z}Z\rho Z.$ (5)
For an error configuration $C$, let $n_{x},n_{y},n_{z}$ denote the numbers of qubits which suffer $X,Y,Z$ errors, respectively. The probability of $C$ is then expressed as
$\operatorname{Prob}(C)$ $\coloneqq$ ${p_{x}}^{n_{x}}\,{p_{y}}^{n_{y}}\,{p_{z}}^{n_{z}}\,(1-p)^{N_{\mathrm{q}}-(n_{x}+n_{y}+n_{z})}$ (6)
$=$ $(1-p)^{N_{\mathrm{q}}}\,\left(\frac{p_{x}}{1-p}\right)^{n_{x}}\,\left(\frac{p_{y}}{1-p}\right)^{n_{y}}\,\left(\frac{p_{z}}{1-p}\right)^{n_{z}}$
$=$ $(1-p)^{N_{\mathrm{q}}}\,\pi(C),$
where
$\pi(C)$ $\coloneqq$ $\left(\frac{p_{x}}{1-p}\right)^{n_{x}}\,\left(\frac{p_{y}}{1-p}\right)^{n_{y}}\,\left(\frac{p_{z}}{1-p}\right)^{n_{z}}$ (7)
$=$ $\left(\frac{p}{1-p}\right)^{\alpha_{x}n_{x}+\alpha_{y}n_{y}+\alpha_{z}n_{z}}$ (8)
is a factor relevant to $p_{x},p_{y},p_{z}$. From Eqs. (7),(8), the coefficients $\alpha_{x},\alpha_{y},\alpha_{z}$ are expressed as
$\alpha_{\mu}=\frac{\log{\frac{p_{\mu}}{1-p}}}{\log{\frac{p}{1-p}}}$ (9)
where $\mu\in\{x,y,z\}$. Assuming that we know or estimate $p_{x}$, $p_{y}$, $p_{z}$ enables optimal decoding, as discussed below.
We define energy of the error configuration $C$ as the weighted sum of the numbers [^2] The numbers $n_{x}$, $n_{y}$, $n_{z}$ are calculated as follows. We formulate and implement the Metropolis method with bit representation taking values in the set $\{0,1\}$, instead of Ising spins taking values in the set $\{1,-1\}$. In an error configuration, an error that suffers each data qubit is represented by two bits. In the two bits, the upper and lower bit means for $X$ and $Z$ error, respectively. Thus, $I$, $Z$, $X$, $Y$ errors are represented as $00$, $01$, $10$, $11$, respectively. As such, counting the numbers of qubits which have bit representations $01$, $10$, $11$ yields $n_{z}$, $n_{x}$, $n_{y}$, respectively. :
$\mathcal{H}(C)\coloneqq\alpha_{x}n_{x}+\alpha_{y}n_{y}+\alpha_{z}n_{z}.$ (10)
Using Eq. (10), Eq. (8) is expressed as
$\pi(C)=\left(\frac{p}{1-p}\right)^{\mathcal{H}(C)}=e^{-\beta\,\mathcal{H}(C)}$ (11)
where $\beta$ is inverse temperature, defined as [^3] The Nishimori temperature includes the coefficient $1/2$ in the formulation by Ising spins [28, 29]. Instead, we employ the formulation by bit representation, and hence, this coefficient does not appear.
$\beta\coloneqq-\log\left(\frac{p}{1-p}\right).$ (12)
Let us define an equivalence class for an error configuration $C$ via the equivalence class $\cdot\,$]. Given a syndrome $S$, we can construct a pure error, i.e., an error configuration $T(S)$ consistent with the syndrome $S$. Using any pure error $T(S)$, any error configuration $C$ consistent with the syndrome $S$ can be decomposed as
$C=T(S)\,L_{P}\,G$ (13)
where $P\in\{I,X,Y,Z\}$ and a stabilizer operator $G\in\mathcal{G}$ and vice versa [30].
Left-multiplying both sides of Eq. (13) by $T(S)$ yields
$T(S)\,C]=[L_{P}].$ (14)
Then, the set of error configurations $C$ satisfying Eq. (14) is defined as
$\operatorname{Rep}_{T(S)}(L_{P})$ $\coloneqq$ $\left\{C\,\middle|\,[T(S)\,C]=[L_{P}]\right\}$ (15)
$=$ $\left\{T(S)\,L_{P}\,G\,\middle|\,G\in\mathcal{G}\right\},$ (16)
which is the set of representatives of another equivalence class, referred to as class for the logical operator $L_{P}$ based on $T(S)$. We employ a pure error $T(S)$ as output of some decoder. Specifically, we utilize our greedy matching decoder, elaborated in Sec. IV.3.
The probability of the class for the logical operator $L_{P}$ based on $T(S)$ is given by
$\mathrm{P}_{L_{P}}\coloneqq\max_{C}\left\{\operatorname{Prob}(C)\,\middle|\,C\in\operatorname{Rep}_{T(S)}(L_{P})\right\}.$ (17)
Eq. (17) is expressed by using Eqs. (6),(11) as
$\mathrm{P}_{L_{P}}\propto e^{-\beta\,E_{L_{P}}}$ (18)
where
$E_{L_{P}}$ $\coloneqq$ $\min_{C}\left\{\mathcal{H}(C)\,\middle|\,C\in\operatorname{Rep}_{T(S)}(L_{P})\right\}$ (19)
$=$ $\min_{G\in\mathcal{G}}\mathcal{H}\left(T(S)\,L_{P}\,G\right)$
which follows from Eq. (16).
We determine the most likely class for the logical operator $L_{P^{*}}$ which maximizes the error probabilities $\mathrm{P}_{L_{P}}$:
$P^{*}\coloneqq\operatorname*{\arg\max}_{P\in\{I,X,Y,Z\}}\mathrm{P}_{L_{P}}=\operatorname*{\arg\min}_{P\in\{I,X,Y,Z\}}E_{L_{P}},$ (20)
which follows from Eq. (18). This decoding principle is based on the theory of the domain-wall of spin-glass models [31, §3.5] [32, 33]. For each $P\in\{I,X,Y,Z\}$, to calculate $E_{L_{P}}$ in Eq. (19), we employ simulated annealing (SA) [21].
For our SA, in the decomposition of Eq. (13), $C_{L_{P}}\coloneqq T(S)\,L_{P}$ is set as an initial configuration of SA, and the stabilizer operator $G$ is updated by the Metropolis method to explore the ground state energy $E_{L_{P}}$. Although it is rigorously proved that SA converges to the optimal solution regardless of the initial configuration if we take an infinite time [34], we expect that in realistic finite time, the initial configuration affects the convergence speed of SA. We examine this effect in Sec. V.3.
We give our SA decoder in Algorithm 1. To acquire the lowest possible energy $E_{L_{P}}$, we carry out simulated annealing multiple times (say $N_{\mathrm{SA}}$ times) for the identical error configuration $C_{L_{P}}$ in Lines 78 of Algorithm 1, and select the minimum energy among them in Line 9. In Line 14, we finally obtain the recovery configuration $C_{L_{P^{*}}}=T\,L_{P^{*}}$. [^4] As a recovery error configuration, any configuration $C\in\operatorname{Rep}_{T(S)}(L_{P^{*}})$ can be selected.
It should be noted that, in the nested for loops in Lines 5 and 7 in Algorithm 1, each iteration is an independent SA task and, hence, these tasks are obviously parallelizable. The total number of iterations in these loops are $4$ and $N_{\mathrm{SA}}$ and, therefore, if parallelized, the elapsed time in parallel execution is expected to be the elapsed time in sequential execution divided by $4N_{\mathrm{SA}}$. This parallelized implementation is left to our future work.
Algorithm 1 invokes the simulated annealing, shown in Algorithm 2. Algorithm 2 performs the Metropolis method, given in Algorithm 3. Line 9 in Algorithm 3 flips spins, corresponding to errors on qubits, by acting a stabilizer generator [^5] These spins are implemented by bits. Hence, the multiplication of error configurations $C\,G_{f}$ is implemented as addition modulo two. in order to keep satisfying the syndrome consistency condition $\partial{C}=S$ throughout the entire SA algorithm.
For the physical error probability $p$, Eq. (12) gives the Nishimori temperature $\beta_{\mathrm{nishimori}}$ where actual error configuration and recovery configuration equally fluctuate, and we can incorporate the degeneracy of error configurations with the same energy. Hence, it is optimal for decoding as a fixed temperature [28, 29, 33].
In Algorithm 2, to accelerate convergence of SA, we employ inverse temperature scheduling with logarithmic function from $\beta_{\mathrm{init}}\coloneqq 0.9\,\beta_{\mathrm{nishimori}}$ to $\beta_{\mathrm{final}}\coloneqq\beta_{\mathrm{nishimori}}$
$\beta_{i}\coloneqq\beta_{\mathrm{init}}\,\left(1+r\,\log{i}\right)\quad\text{for $i=1,\dots,N_{\mathrm{\beta}}$}$ (21)
with
$r\coloneqq\left.\left(\dfrac{\beta_{\mathrm{final}}}{\beta_{\mathrm{init}}}-1\right)\middle/\left(\log{N_{\mathrm{\beta}}}\right)\right.$. (22)
For each inverse temperature $\beta$, we execute the Metropolis step $N_{\mathrm{stabilizer}}$ times, which is counted as one Monte Carlo step, according to the convention.
Algorithm 1 Decoder using simulated annealing
1: $S$ : Syndrome for error configuration
2: Recovery configuration $C$
3: function SA_Decoder($S$)
4:   $T\leftarrow$ greedy_matching($S$) $\triangleright$ pure error
5:   for all $P\in\{I,X,Z,Y\}$ do
6:    $C_{L_{P}}\coloneqq T\,L_{P}$ $\triangleright$ Initial configuration for SA
7:    for $i=1$ to $N_{\mathrm{SA}}$ do
8:      $E^{(i)}_{L_{P}}\leftarrow$ simulated_annealing($C_{L_{P}}$)
9:    $E_{L_{P}}\coloneqq\min\limits_{i=1,\dots,N_{\mathrm{SA}}}{E^{(i)}_{L_{P}}}$
10:
11:
12:   $P^{*}\coloneqq\operatorname*{\arg\min}\limits_{P\in\{I,X,Y,Z\}}{E_{L_{P}}}$ $\triangleright$ Eq. (20) 13:
14:   return $C_{L_{P^{*}}}$
Algorithm 2 Simulated annealing
1: $C$ : Error configuration
2: Ground state energy (or lower excited state energy close to the ground state energy)
3: function Simulated_Annealing($C$)
4:   $E\leftarrow\mathcal{H}(C)$ $\triangleright$ Eq. (10)
5:   $E_{\mathrm{min}}\leftarrow E$
6:   $r\coloneqq\left.\left(\dfrac{\beta_{\mathrm{final}}}{\beta_{\mathrm{init}}}-1\right)\middle/\left(\log{N_{\mathrm{\beta}}}\right)\right$. $\triangleright$ Eq. (22) 7:
8:   for $i=1$ to $N_{\mathrm{\beta}}$ do
9:    $\beta\leftarrow\beta_{\mathrm{init}}\,\left(1+r\,\log{i}\right)$ $\triangleright$ Eq. (21)
10:    for $N_{\mathrm{stabilizer}}$ times do
11:      $\Delta E\leftarrow$ metropolis_Step($\beta$,  $C$)
12:      $E\leftarrow E+\Delta E$
13:      $E_{\mathrm{min}}\leftarrow\min\{E_{\mathrm{min}},\,E\}$ $\triangleright$ Update min energy
14:
15:   return $E_{\mathrm{min}}$
Algorithm 3 Metropolis method which acts a stabilizer generator
1: $\beta$ : inverse temperature,  $C$ : Error configuration
2: Difference of energy of error configurations
3: function Metropolis_Step($\beta$,  $C$)
4:   Select a stabilizer generator $G_{f}$ randomly.
5:   $\Delta E\leftarrow\mathcal{H}(C\,G_{f})-\mathcal{H}(C)$ 6:
7:   Generate a uniform random probability $r\in[0,1]$.
8:   if $\Delta E<0$ or $r<e^{-\beta\,\Delta E}$ then
9:    $C\leftarrow C\,G_{f}$ $\triangleright$ Flips spins
10:    return $\Delta E$
11:   else
12:    return 0
IV.2 The MWPM decoder
IV.2.1 The MWPM decoding algorithm
We briefly review the algorithm of the MWPM decoder for the XZZX code [9].
The MWPM decoder is formulated as an algorithm on the decoding graph $(\mathcal{V}_{\bullet},\mathcal{E}_{\bullet})$, introduced in Sec. III. For $u,v\in S_{\bullet}$, we define the weighted distance $w(u,v)$ between the vertices $u=(u_{x},u_{y})$ and $v=(v_{x},v_{y})$ as the shorter of the following two path lengths:
  • The weighted Manhattan distance
    $\alpha_{z}\lvert u_{x}-v_{x}\rvert\,+\,\alpha_{x}\lvert u_{y}-v_{y}\rvert$, (23)
  • The length of ‘path via a virtual boundary vertex’ [35]. For $u,v\in S_{\mathrm{gray}}$, this path consists of only horizontal edges, which connect $u$ and the left or right boundary and connect $v$ and the left or right boundary. The length of the path equals to
[…]
    For $u,v\in S_{\mathrm{white}}$, this path consists of only vertical edges, which connect $u$ and the top or bottom boundary and connect $v$ and the top or bottom boundary. The length of the path equals to
[…]
We present the algorithm of the MWPM decoder in Algorithm 4. In Line 4, if the number of vertices $|S|$ is odd, to enable a perfect matching, the virtual boundary vertex is appended to the syndrome $S$ so that $|S|$ is an even number [35]. In Line 6, we create the list of all possible candidates for pairing
$K\coloneqq\{\text{unordered pair }(u,v)\,|\,u,v\in S_{\bullet},\,u\neq v\}.$ (26)
In Line 8, the MWPM_graph function finds a minimum-weight perfect matching (MWPM) [36], i.e., a perfect matching $L\subset K$ such that the total distance $\sum_{(u,v)\in L}w(u,v)$ is minimum. For our implementation, we employ the C++ library, Blossom V [37].
Algorithm 4 The MWPM decoder
1: $S_{\bullet}$ : Syndrome with $\bullet\in\{\mathrm{white},\mathrm{gray}\}$
2: Recovery chain $C_{\bullet}$ consistent with $S_{\bullet}$
3: function MWPM_chain($S_{\bullet}$)
4:   if $|S_{\bullet}|$ is odd then
5:    Append ‘the virtual boundary vertex’ to $S_{\bullet}$
6:   $K\coloneqq\{\text{unordered pair }(u,v)\,|\,u,v\in S_{\bullet},\,u\neq v\}$ $\triangleright$ Eq. (26) 7:
8:   $L\leftarrow$ MWPM_graph($K$) $\triangleright$ Utilize Blossom V 9:
10:   $C_{\bullet}\leftarrow\emptyset$
11:   for all $(u,v)\in L$ do
12:    Add edges which connect $u$ and $v$ to $C_{\bullet}$.
13:
14:   return $C_{\bullet}$
1: $S\coloneqq(S_{\mathrm{white}},\,S_{\mathrm{gray}})$ : Syndrome
2: Recovery configuration $C$ consistent with $S$
3: function MWPM($S$)
4:   $C_{\mathrm{white}}\leftarrow$ MWPM_chain($S_{\mathrm{white}}$)
5:   $C_{\mathrm{gray}}\leftarrow$ MWPM_chain($S_{\mathrm{gray}}$)
6:   $C\coloneqq(C_{\mathrm{white}},C_{\mathrm{gray}})$ 7:
8:   return $C$
In Line 12, for each matching pair $(u,v)\in L$, we connect the vertices $u$ and $v$ with a path which consists of edges, i.e., applying $Z$ ($X$) operators to the qubits located at the centers of horizontal (vertical) edges, respectively. The set of these edges forms a recovery chain $C_{\bullet}$.
The MWPM function invokes the MWPM_CHAIN function twice, i.e., for the syndromes $S_{\mathrm{white}}$ and $S_{\mathrm{gray}}$ on the white gray and faces, respectively. Note that, the syndromes on only faces of the same color may be connected.
The MWPM algorithm is too complicated to implement in a hardware system such as FPGA. Thus, we propose a simple variant algorithm in Sec. IV.3.
IV.2.2 Interpreting the MWPM decoder as an energy minimization problem
Let us interpret the MWPM decoder as energy minimization problem, then compare it with our SA decoder.
Since the perfect matching is equivalent to the syndrome consistency $\partial{C_{\bullet}}=S$, the total distance of such an error chain $C_{\bullet}$ is defined as
$w(C_{\bullet})\coloneqq\sum_{(u,v)\in L_{0}}w(u,v)$ (27)
where $L_{0}$ is a perfect matching. Then, the MWPM decoding algorithm is equivalent to find a recovery error chain $C_{\bullet}$:
$\arg\min_{C_{\bullet}}\left\{w(C_{\bullet})\,\middle|\,\partial{C_{\bullet}}=S\right\}.$ (28)
Since the weighted distance $w(u,v)$ is defined by Eqs. (23),(24) or (25), Eq. (27) is expressed as
$w(C_{\bullet})=\alpha_{z}\,n_{\mathrm{hor}}\,+\,\alpha_{x}\,n_{\mathrm{ver}}$ (29)
where $n_{\mathrm{hor}}$ ($n_{\mathrm{ver}}$) is the number of errors which reside on horizontal (vertical) edges in the error chain $C_{\bullet}$, respectively. In Eq. (29), the weight coefficients $\alpha_{x},\alpha_{z}$ can be calculated by Eq. (9) from the error probabilities $p_{x},p_{z}$ assuming we know.
Let us compare the total distance in Eq. (29) and the energy defined in Eq. (10). Since Eq. (29) does not contain the probability $p_{y}$, the MWPM decoder cannot properly consider $Y$ errors unless $p_{x}:p_{y}:p_{z}=1:2:1$, and therefore is not optimal. On the other hand, Eq. (10) contains the probability $p_{y}$, and hence our SA decoder is able to attain optimal accuracy. In Sec. V, we execute the MWPM decoder to evaluate the accuracy of our SA decoders for comparison.
IV.3 The greedy matching decoder
As described in Sec. IV.2.1, the MWPM decoder utilizes the MWPM graph algorithm, i.e., the perfect matching giving the minimum of the total distance. Exchanging this ‘global’ minimality for a low computational cost and simple algorithm yields a greedy version of the perfect matching algorithm. In this subsection, we elaborate on our greedy matching algorithm. We employ this algorithm to construct a pure error $T(S)$, set as an initial configuration for our SA decoder in Algorithm 1.
We give our algorithm in Algorithm 5. In Line 10, the GREEDY_MATCHING_CHAIN function matches the closest pair $(u,v)\in K$ with respect to $w(u,v)$, and connects the vertices $u$ and $v$. In Line 12, it excludes all the pairs whose vertex is $u$ or $v$ from the candidate list $K$. This algorithm thus enables us to monotonically reduce the size of the list $K$.
We efficiently implement $K$ using priority queue [38, §6.5]. In the priority queue with $n$ elements (e.g., a binary heap), each of insertion and extract-min operations takes $\mathcal{O}(\log n)$; therefore, performing the operations for $n$ elements has time complexity $\mathcal{O}(n\log n)$. For our greedy matching decoder, the size of the priority queue is $n=\mathcal{O}(|S_{\bullet}|^{2})$, where $|S_{\bullet}|$ denotes the number of incorrect syndromes. Thus, the time complexity of our greedy matching decoder is $\mathcal{O}(|S_{\bullet}|^{2}\log{|S_{\bullet}|})$.
Algorithm 5 The greedy matching decoder
1: $S_{\bullet}$ : Syndrome with $\bullet\in\{\mathrm{white},\mathrm{gray}\}$
2: Recovery chain $C_{\bullet}$ consistent with $S_{\bullet}$
3: function greedy_matching_chain($S_{\bullet}$)
4:   if $|S_{\bullet}|$ is odd then
5:    Append ‘the virtual boundary vertex’ to $S_{\bullet}$
6:   $K\coloneqq\{\text{unordered pair }(u,v)\,|\,u,v\in S_{\bullet},\,u\neq v\}$ $\triangleright$ Eq. (26)
7:   $C_{\bullet}\leftarrow\emptyset$ 8:
9:   while $K$ is not empty do
10:    Pick a pair $(u,v)\in K$ such that $w(u,v)$ is minimum.
11:    Add edges which connect $u$ and $v$ to $C_{\bullet}$.
12:    Exclude all the pairs whose vertex is $u$ or $v$ from the list $K$.
13:
14:   return $C_{\bullet}$
1: $S\coloneqq(S_{\mathrm{white}},\,S_{\mathrm{gray}})$ : Syndrome
2: Recovery configuration $C$ consistent with $S$
3: function greedy_matching($S$)
4:   $C_{\mathrm{white}}\leftarrow$ greedy_matching_chain($S_{\mathrm{white}}$)
5:   $C_{\mathrm{gray}}\leftarrow$ greedy_matching_chain($S_{\mathrm{gray}}$)
6:   $C\coloneqq(C_{\mathrm{white}},C_{\mathrm{gray}})$ 7:
8:   return $C$
As well as the MWPM decoder, since our greedy matching decoder employs the weighted distance $w(u,v)$ which is calculated from the probabilities $p_{z}$ and $p_{x}$, it cannot take into account the probability $p_{y}$.
IV.4 Our SA decoder which starts with different pure errors
In Sec. IV.1, we described our SA decoder in Algorithm 1, which performs SA multiple times to achieve energy as low as possible.
To improve efficiency, rather than using an identical pure error repeatedly, it is preferable to use a different pure error for each SA run. To this end, we modify our greedy matching algorithm. Specifically, when multiple pairs of vertices have the same distance, their matching order is randomized using pseudo-random numbers. This modified algorithm is presented in Algorithm 6. In Algorithm 6, we create the candidate list $K_{w}$, defined as the set of pairs which give the distance $w$:
$K_{w}\coloneqq\{(u,v)\in K\,|\,w(u,v)=w\}.$ (30)
This newly introduced randomness may produce a different pure error for every invocation, while still remaining consistent with the given syndrome $S$. However, such pure errors differ not only in their stabilizer components but also in their logical components in the decomposition of Eq. (13). Therefore, we need to transform a logical operator to the one based on a ‘standard’ pure error, as follows.
We present SA decoder which utilizes Algorithm 6 in Algorithm 7. In Line 5, for the first SA, we construct a pure error $T_{1}(S)$, which is used as ‘standard’.
Let us denote by $T(S)$ a pure error constructed from the second SA onward. By applying Eq. (13), $T_{1}(S)$ can be decomposed as
$T_{1}(S)=T(S)\,L_{Q}\,G^{\prime}$ (31)
for $Q\in\{I,X,Z,Y\}$ and $G^{\prime}\in\mathcal{G}$. The logical operator $L_{Q}$ expresses ‘difference’ of equivalence classes of logical operators based on $T(S)$ and $T_{1}(S)$, and is obtained by $L_{Q}]=[T(S)\,T_{1}(S)]$.
From Eq. (31), it follows that
$\operatorname{Rep}_{T(S)}(L_{P})=\operatorname{Rep}_{T_{1}(S)}(L_{QP}),$
which implies that “the logical operator $L_{P}$ based on T(S)T(S)” equals to “the logical operator $L_{QP}$ based on $T_{1}(S)$”. Thus, for each $P\in\{I,X,Z,Y\}$, we store the calculated energy in $E^{(i)}_{L_{QP}}$ in Line 16.
It should be noted that, in the for loops in Lines 11 and 14 in Algorithm 7, each iteration is an independent task, and hence these tasks can obviously be parallelized. In addition, Line 13 can be parallelized with Lines 1516.
Algorithm 6 The greedy matching decoder which may produce a different pure error
1: $S_{\bullet}$ : Syndrome with $\bullet\in\{\mathrm{white},\mathrm{gray}\}$
2: Recovery chain $C_{\bullet}$ consistent with $S_{\bullet}$
3: function greedy_matching_diff_chain($S_{\bullet}$)
4:   if $|S_{\bullet}|$ is odd then
5:    Append ‘the virtual boundary vertex’ to $S_{\bullet}$
6:   $K\coloneqq\{\text{unordered pair }(u,v)\,|\,u,v\in S_{\bullet},\,u\neq v\}$ $\triangleright$ Eq. (26)
7:
$C_{\bullet}\leftarrow\emptyset$
8:
9:   while $K$ is not empty do
10:    $w\coloneqq\min\{w(u,v)\,|\,(u,v)\in K\}$
11:    $K_{w}\coloneqq\{(u,v)\in K\,|\,w(u,v)=w\}$ $\triangleright$ Eq. (30)
12:    while $K_{w}$ is not empty do
13:      Pick a pair $(u,v)\in K_{w}$ randomly.
14:      Add edges which connect $u$ and $v$ to $C_{\bullet}$.
15:      Exclude all the pairs whose vertex is $u$ or $v$ from the lists $K_{w}$ and $K$.
16:
17:   return $C_{\bullet}$
1: $S\coloneqq(S_{\mathrm{white}},\,S_{\mathrm{gray}})$ : Syndrome
2: Recovery configuration $C$ consistent with $S$
3: function greedy_matching_different($S$)
4:   $C_{\mathrm{white}}\leftarrow$ greedy_matching_diff_chain($S_{\mathrm{white}}$)
5:   $C_{\mathrm{gray}}\leftarrow$ greedy_matching_diff_chain($S_{\mathrm{gray}}$)
6:   $C\coloneqq(C_{\mathrm{white}},C_{\mathrm{gray}})$ 7:
8:   return $C$
Algorithm 7 Decoder using simulated annealing which may start with different pure errors
1: $S$ : Syndrome for error configuration
2: Recovery configuration $C$
3: function SA_Decoder_Different_pure_error($S$)
4:   $\triangleright$  First SA   $\triangleleft$
5:   $T_{1}\leftarrow$ greedy_matching_different($S$)
6:   for all $P\in\{I,X,Z,Y\}$ do
7:    $C^{(1)}_{L_{P}}\coloneqq T_{1}\,L_{P}$ $\triangleright$ Initial configuration for SA
[…]
9:
10:   $\triangleright$  Second and later SAs   $\triangleleft$
11:   for $i=2$ to $N_{\mathrm{SA}}$ do
12:    $T\leftarrow$ greedy_matching_different($S$)
13:    Determine $Q\in\{I,X,Z,Y\}$ such that $L_{Q}]=[T\,T_{1}]$.
14:    for all $P\in\{I,X,Z,Y\}$ do
15:      $C_{L_{P}}\coloneqq T\,L_{\mathrm{P}}$ $\triangleright$ Initial configuration for SA
16:      $E^{(i)}_{L_{QP}}\leftarrow$ simulated_annealing($C_{L_{P}}$)
17:
18:   for all $P\in\{I,X,Z,Y\}$ do
19:    $E_{L_{P}}\coloneqq\min\limits_{i=1,\dots,N_{\mathrm{SA}}}{E^{(i)}_{L_{P}}}$
20:
21:   $P^{*}\coloneqq\operatorname*{\arg\min}\limits_{P\in\{I,X,Y,Z\}}{E_{L_{P}}}$ $\triangleright$ Eq. (20) 22: 23:
24:   return $C^{(1)}_{L_{P^{*}}}$
Like the MWPM decoder, our greedy matching decoder is MLE-type; the MLD variant is obtained by disabling SA (i.e., by setting $N_{\beta}=0$) in Algorithm 7.
V Numerical result
We perform Monte Carlo simulations of our SA decoder under the setting described in Sec. V.1. In Sec. V.2, we evaluate the logical error rate of our SA decoder. In Sec. V.3, we examine the validity of our greedy matching decoder used to construct initial configurations. We compare elapsed times of our SA decoder with other decoding algorithms in Sec. V.4.
V.1 Simulation setting
In the Monte Carlo simulation, we generate samples of actual error configurations $C$ by means of pseudo-random numbers, drawn from the physical error rate $p$ and the error bias $p_{x}:p_{y}:p_{z}$ we specify. Performing the stabilizer measurement on the actual error configurations $C$ yields the syndrome $S$. Then, we execute each decoder to obtain the recovery configuration $\bar{C}$. If the product $C\,\bar{C}$ belongs to the (non-trivial) logical operator class $L_{X}],[L_{Y}]$ or $L_{Z}]$, then the logical error occurs. By repeating the aforementioned procedure $N_{\mathrm{sample}}$ times, we estimate the logical error rate $P_{\mathrm{L}}$. Except in Sec. V.4, we employ $N_{\mathrm{sample}}=10^{6}$ to demand sufficient statistical precision. Except in Sec. V.3, we refer to Algorithm 7 as ‘our SA decoder’ and employ the number of inverse temperatures $N_{\beta}=100$.
Although our SA decoder can be applied to any type of biased noise, we focus on $Y$-biased noise such that $X$ and $Z$ errors occur with the same probabilities $p_{x}=p_{z}$, which is the simplest case where the MWPM decoder cannot produce a correct recovery configuration. By the condition $p_{x}=p_{z}$, we can express as $p_{x}:p_{y}:p_{z}=1:\eta:1$ where $\eta$ indicates the $Y$-biased factor. Since the depolarizing noise is the case where $p_{x}=p_{y}=p_{z}=1/3$, it corresponds to $p_{x}:p_{y}:p_{z}=1:1:1$, i.e., $\eta=1$. The more $Y$-biased, the less accurate the MWPM decoder.
V.2 Logical error rates
Let us compare the decoding accuracy of our SA decoder described in Algorithm 7 and the CPLEX decoder described in Sec. A. First, we examine how the logical error rate varies with the code distance $d$.
Fig. 3 depicts the logical error rates $P_{\mathrm{L}}$ under the depolarizing noise and for the code distances $d=5,7,9$. In Fig. 3, for every $d$, our SA decoder achieves CPLEX’s rates, which is optimal among MLE-type decoders. In lower physical error rates $p$, our SA decoder outperforms the CPLEX decoder. This is because, unlike the CPLEX decoder, our SA decoder accounts for the degeneracy of error configurations by setting the final temperature to the Nishimori temperature $\beta_{\mathrm{nishimori}}$ in Eq. (12).
For both the SA and CPLEX decoders, from the intersection point of these curves, we determine the threshold $p_{\mathrm{th}}\simeq 0.18$, which approximately equal to $p_{\mathrm{th}}=0.187$ obtained by the MPS decoder [9]. Below the threshold $p_{\mathrm{th}}$, as $d$ increases, the logical error rate decreases, as expected.
Figure 3: Logical error rates under the depolarizing noise and the code distance $d=5,7,9$. Our SA decoder achieves the logical error rates of the CPLEX decoder for all $d$. We employ $N_{\mathrm{SA}}=100$, which suffices to attain CPLEX’s accuracy. The error bars indicate one standard deviation.
Fig. 4 shows the logical error rates under the $Y$-biased noise where $p_{x}:p_{y}:p_{z}=1:5:1$. In Fig. 4, for every $d$, our SA decoder achieves CPLEX’s rates, which are optimal. For both SA and CPLEX decoders, the intersection point of these curves lies around $p=0.2$, which we consider to be the threshold. Below this threshold, the logical error rate decreases as $d$ increases.
Figure 4: Logical error rates under the $Y$-biased noise where $p_{x}:p_{y}:p_{z}=1:5:1$ and for the code distance $d=5,7,9$. Our SA decoder achieves the logical error rates of the CPLEX decoder. We employ $N_{\mathrm{SA}}=100$, which suffices to attain CPLEX’s accuracy. The error bars indicate one standard deviation.
Fig. 5 shows the logical error rates under the $Y$-biased noise where $p_{x}:p_{y}:p_{z}=1:10:1$. In Fig. 5, no intersection point is observed for $d=5,7,9$; hence, we infer that the threshold lies beyond the plotted range, i.e., $p_{\mathrm{th}}>0.2$. On the other hand, for $d=9$ in Fig. 5, our SA decoder does not achieve CPLEX’s rate. The curves for $d=5,7,9$ do not intersect, due to the critical slowing down of the MCMC method. For our SA decoder, the curves for $d=7,9$ intersect around $p=0.11$, which we regarded as the threshold. When $p<0.11$, the logical error rate decreases as $d$ increases.
Figure 5: Logical error rates under the $Y$-biased noise where $p_{x}:p_{y}:p_{z}=1:10:1$ and for the code distance $d=5,7,9$. Our SA decoder cannot achieve the logical error rate of the CPLEX decoder for $d=9$, even by $N_{\mathrm{SA}}=100$. The error bars indicate one standard deviation.
Next, we investigate the dependence of the logical error rates $P_{\mathrm{L}}$ on the number of SA, $N_{\mathrm{SA}}$ for fixed code distance $d=9$. For comparison, we also calculate the rates by our greedy matching decoder, i.e., the GREEDY_MATCHING function described in Algorithm 5 and the MWPM decoder i.e., the MWPM function described in Algorithm 4.
Fig. 6 shows the logical error rates under the depolarizing noise. Figs. 7 and 8 show the logical error rates under $Y$-biased noises where $p_{x}:p_{y}:p_{z}=1:5:1$ and $1:10:1$ for several values of $N_{\mathrm{SA}}$, respectively.
Figure 6: Logical error rates under the depolarizing noise and $N_{\mathrm{SA}}=1,\dots,10$. We employ the code distance $d=9$. The error bars indicate one standard deviation.
Figure 7: Logical error rates under $Y$-biased noise where $p_{x}:p_{y}:p_{z}=1:5:1$ and $N_{\mathrm{SA}}=1,\dots,100$. We employ the code distance $d=9$. The error bars indicate one standard deviation.
Figure 8: Logical error rates under $Y$-biased noise where $p_{x}:p_{y}:p_{z}=1:10:1$ and $N_{\mathrm{SA}}=1,\dots,100$. We employ the code distance $d=9$. The error bars indicate one standard deviation.
In Figs. 68, we observe that the logical error rate $P_{\mathrm{L}}$ of our SA decoder monotonically decreases as $N_{\mathrm{SA}}$ increases. For $N_{\mathrm{SA}}\geq 2$, our SA decoder achieves logical error rates lower than our greedy matching decoder and is therefore an improvement over our greedy matching. For $N_{\mathrm{SA}}\geq 3$, our SA decoder is more accurate than the MWPM decoder.
To achieve the accuracy by the CPLEX decoder, under the depolarizing noise in Fig. 6, $N_{\mathrm{SA}}=10$ is sufficient. Although under the noise $p_{x}:p_{y}:p_{z}=1:5:1$ in Fig. 7, $N_{\mathrm{SA}}=100$ is almost sufficient. Under the noise $p_{x}:p_{y}:p_{z}=1:10:1$ in Fig. 8, $N_{\mathrm{SA}}=100$ is not sufficient. Hence, to attain the optimal logical error rate, we observe that the more strongly $Y$-biased noise, the more $N_{\mathrm{SA}}$ our SA decoders need.
V.3 Convergence of the SA decoders
To construct a pure error, there are the following three methods:
Boundary
Every incorrect syndrome is directly connected with closer one in the top and bottom boundaries for white faces (the left and right boundaries for gray faces). We employ the SA decoder described in Algorithm 1.
Greedy matching (same)
It is Algorithm 5 invoked in the SA decoder of Algorithm 1, described in Sec. IV.3. For every invocation, it produces the same pure error.
Greedy matching (different)
It is Algorithm 6 invoked in the SA decoder of Algorithm 7, described in Sec. IV.4. For every invocation, it may produce a different pure error. We use this method in the present paper, unless otherwise stated.
To compare the effect of these methods on the convergence of our SA decoder, we evaluate the ratio of logical error rates
$R_{\mathrm{L}}\coloneqq\left(\dfrac{P^{\mathrm{(SA)}}_{\mathrm{L}}}{P^{\mathrm{(CPLEX)}}_{\mathrm{L}}}\right)^{\frac{2}{d+1}}$ (32)
where $P^{\mathrm{(SA)}}_{\mathrm{L}}$ and $P^{\mathrm{(CPLEX)}}_{\mathrm{L}}$ are the logical error rates obtained by the SA and CPLEX decoders, respectively. Here, the power $2/(d+1)$ is to cancel out $d$-dependence. As for the ratio $R_{\mathrm{L}}$, the lower value implies that our SA decoder is more accurate than the CPLEX decoder, and the value $1$ indicates that our SA decoder attains CPLEX decoder’s accuracy.
Fig. 9 shows the ratio $R_{\mathrm{L}}$ for varying the number of inverse temperatures $N_{\beta}$ under the depolarizing noise. Similarly, Fig. 10 and Fig. 11 show the ratio $R_{\mathrm{L}}$ under the $Y$-biased noises where $p_{x}:p_{y}:p_{z}=1:5:1$ and $1:10:1$, respectively. In Figs. 911, the ratio values $R_{\mathrm{L}}$ obtained by the greedy matching and MWPM decoders are drawn by dotted gray and dashed-dotted brown horizontal lines, respectively.
(a) Refer to caption(b) Refer to caption (c) Refer to caption
Figure 9: Convergence of our SA decoder for different construction methods of pure errors under the depolarizing noise. The horizontal axis is the number of inverse temperatures $N_{\beta}$, and the vertical axis is the ratio $R_{\mathrm{L}}$. We employ $d=9$ and $N_{\mathrm{SA}}=10$. (a) $p=0.04$ (b) $p=0.1$ (c) $p=0.15$.
(a) Refer to caption(b) Refer to caption (c) Refer to caption
Figure 10: Convergence of our SA decoder for different construction methods of pure errors under the $Y$-biased noise such that $p_{x}:p_{y}:p_{z}=1:5:1$. The horizontal axis is the number of inverse temperatures $N_{\beta}$, and the vertical axis is the ratio $R_{\mathrm{L}}$. We employ $d=9$ and $N_{\mathrm{SA}}=10$. (a) $p=0.04$ (b) $p=0.1$ (c) $p=0.15$.
(a) Refer to caption(b) Refer to caption (c) Refer to caption
Figure 11: Convergence of our SA decoder for different construction methods of pure errors under the $Y$-biased noise such that $p_{x}:p_{y}:p_{z}=1:10:1$. The horizontal axis is the number of inverse temperatures $N_{\beta}$, and the vertical axis is the ratio $R_{\mathrm{L}}$. We employ $d=9$ and $N_{\mathrm{SA}}=10$. (a) $p=0.04$ (b) $p=0.1$ (c) $p=0.15$.
In Figs. 911, under all the noises and for all the pure error construction methods, we can see that the ratio $R_{\mathrm{L}}$ broadly decreases to the value $1$ as $N_{\beta}$ increases. For $N_{\beta}\geq 20$, the ‘Greedy matching (same)’ and ‘Greedy matching (different)’ methods outperform the MWPM decoder.
Setting $N_{\beta}=0$ disables SA. Hence, $N_{\mathrm{SA}}$ indicates the number of not SAs but pure errors. In this setting, the ‘Greedy matching (same)’ and ‘Greedy matching (different)’ decoders become their MLD variants. In all the cases (a), (b), (c) of Figs. 911, the ‘Greedy matching (different)’ method with $N_{\beta}=0$ achieves higher accuracy than the MWPM decoder.
For all values of $N_{\beta}$ in Figs. 911, ‘Greedy matching (different)’ is the highest accurate, followed by ‘Greedy matching (same)’, while the ‘Boundary’ method is the least accurate. The causes of these accuracy differences are considered as follows:
  • Whereas the ‘Boundary’ method ignores the actual error probabilities $p_{x}$, $p_{y}$, $p_{z}$, the ‘Greedy matching (same)’ and ‘Greedy matching (different)’ methods take $p_{x}$, $p_{z}$ (but not $p_{y}$) into account. Consequently, the initial error configuration produced by ‘Greedy matching (same)’ is typically more similar to the actual error configuration than that produced by the ‘Boundary’ method. Thus, the ‘Greedy matching (same)’ method is more accurate than the ‘Boundary’ method.
  • The ‘Greedy matching (different)’ method may produce a different initial error configuration for each SA. It leads to numerous variations of initial configurations, increasing the chance of reaching lower-energy states and thereby reducing the logical error rate.
V.4 Comparison of decoding time with other decoding algorithms
We now evaluate the decoding time of our SA decoder.
First, Fig. 12 shows the elapsed times for the SA and CPLEX decoders for various code distances $d$ under the depolarizing noise. Similarly, Figs. 13 and 14 show the elapsed times under $Y$-biased noises where $p_{x}:p_{y}:p_{z}=1:5:1$ and $1:10:1$, respectively. For measuring the elapsed times, we execute the SA and CPLEX decoders without any parallelization. From Figs. 1214, we can see that the larger $d$, the longer time both the SA and CPLEX decoders take.
(a) Refer to caption (b) Refer to caption (c) Refer to caption
Figure 12: Elapsed times for the SA and CPLEX decoders under the depolarizing noise. The vertical axis is logarithmic scale. (a) $p=0.02$ (b) $p=0.1$ (c) $p=0.15$. In (a) and (b), the dashed orange line implies elapsed time divided by $4N_{\mathrm{SA}}$, which is time assuming parallel execution of every SA. In (c), the CPLEX decoder indicates exponential time for the code distance $d$.
(a) Refer to caption (b) Refer to caption (c) Refer to caption
Figure 13: Elapsed times for the SA and CPLEX decoders under the $Y$-biased noise where $p_{x}:p_{y}:p_{z}=1:5:1$. The vertical axis is logarithmic scale. (a) $p=0.02$ (b) $p=0.1$ (c) $p=0.15$. In (a) and (b), the dashed orange line implies elapsed time divided by $4N_{\mathrm{SA}}$, which is time assuming parallel execution of every SA. In (c), the CPLEX decoder indicates exponential time for the code distance $d$.
(a) Refer to caption (b) Refer to caption (c) Refer to caption
Figure 14: Elapsed times for the SA and CPLEX decoders under the $Y$-biased noise where $p_{x}:p_{y}:p_{z}=1:10:1$. The vertical axis is logarithmic scale. (a) $p=0.02$ (b) $p=0.1$ (c) $p=0.15$. In (a) and (b), the dashed orange line implies elapsed time divided by $4N_{\mathrm{SA}}$, which is anticipated time assuming parallel execution of every SA. In (c), the CPLEX decoder indicates exponential time for the code distance $d$.
For each decoder, we examine the dependence on $d$ and $p$. For all the physical error rates $p$, our SA decoder requires the same time, because the computational cost is determined solely by $N_{\beta}$ and $d$. Recalling that for each inverse temperature, we perform the Metropolis step $\mathcal{O}(N_{\mathrm{stabilizer}})=\mathcal{O}(d^{2})$ times, the total elapsed time scales accordingly. On the other hand, the CPLEX decoder exhibits exponential scaling in computation time with respect to the code distance $d$. This exponentiality is clearly visible for the case (c) $p=0.15$ in Figs. 1214, where it is prohibitive to calculate for $d>11$.
Let us compare the elapsed times of our SA decoder and the CPLEX decoder. For $p=0.15$, our SA decoder is clearly faster than the CPLEX decoder, whereas for $p=0.02,0.1$, it requires more time than the CPLEX decoder.
In each of Figs. 1214, ‘SA (divided by $4N_{\mathrm{SA}}$)’ drawn with the dashed orange line is the elapsed time divided by $4N_{\mathrm{SA}}=40$, indicating anticipated time in parallelized execution. By comparing the dashed orange line and the blue line, we can see that our SA decoder is faster than the CPLEX decoder. Consequently, if our SA decoder is parallelized, it outperforms the CPLEX decoder for all the values $d$ and $p$ we employ.
Next, we present the elapsed times for various decoding algorithms in Fig. 15.
[…]
For our SA decoder, we set the number of inverse temperatures $N_{\mathrm{\beta}}=100$ and $N_{\mathrm{SA}}=10$ in Algorithm 7. Note that $N_{\mathrm{SA}}$ can be any number, because every SAs have almost the same computational cost. For the MPS decoder, we used the implementation for the CSS surface code in the repository [39], because the decoder for the XZZX code employed in Ref. [9] is not open to the public. Here, we employ the truncation dimension $\chi=8$ for the MPS decoder.
From Fig. 15, we observe that our SA decoder is faster than the MPS decoder except for $d=9$. This is because our SA decoder is implemented in C++, whereas the MPS decoder is implemented in Python. As mentioned in Ref. [17, 8], the MPS decoder has a computational complexity of $\mathcal{O}(N_{q})=\mathcal{O}(d^{2})$, which is the same as our SA decoder.
Our SA decoder labeled by ‘SA’ is slower than the CPLEX decoder. However, it can be readily parallelized. Thus, we can consider the elapsed time divided by $4N_{\mathrm{SA}}$, denoted by ‘SA (divided by $4N_{\mathrm{SA}}$)’, again, which takes less time than ‘CPLEX’. Consequently, among the decoders which can deal with the error probability $p_{y}$, such as the SA, CPLEX and MPS decoders, our SA decoder is the fastest if parallelized.
From Fig. 15, we observe that our greedy matching decoder is faster than the MWPM decoder. This result is consistent with the fact that the worst-case computational complexity of our greedy matching decoder is $\mathcal{O}(|S_{\bullet}|^{2}\log{|S_{\bullet}|^{2}})=\mathcal{O}(d^{4}\log{d})$, as discussed in Sec. IV.3, which is lower than the worst-case complexity of the MWPM decoder $\mathcal{O}(|S_{\bullet}|^{3}\log{|S_{\bullet}|})=\mathcal{O}(d^{6}\log{d})$ in Ref. [37]. Note that whereas in circuit-level noise model, the MWPM decoder uses the Dijkstra algorithm to find a path with minimum weight between any possible pair of incorrect syndromes, in the code capacity noise model, it does not because the weights are explicitly given by Eqs. (23),(24) or (25).
Note that the MWPM decoder has an average computational complexity of $\mathcal{O}(d^{2})$ from numerical simulation [35]. Investigating the average complexity for our greedy matching decoder would be meaningful, and is left to our future work.
VI Discussion
As in Sec. V.3, under much more strongly $Y$-biased noise, our SA decoder requires a large $N_{\mathrm{SA}}$ in order to attain the optimal logical error rate. A promising remedy to this problem is to modify the definition of stabilizer generators to include $Y$ operators, analogously to the tailored code for the CSS surface code [8]. Specifically, for the case $p_{y}\geq p_{z}\geq p_{x}$, we should employ ‘XYYX code’. This modified code can detect the endpoints of error chains which consist of $Z$ ($Y$) errors on horizontal (vertical) edges, respectively. Thus, in our greedy matching algorithm, we can take into account the probabilities $p_{z}$ and $p_{y}$. We therefore expect to obtain a pure error, closely resemblant to the actual recovery configuration. By employing this ‘good’ initial configuration, our SA decoder is expected to quickly converge to the optimal logical error rate. Such a treatment can be applied to accommodate the given noise bias.
A key advantage of our SA decoder is that it provides explicit control over the trade-off between computational cost and accuracy by adjusting $N_{\beta}$. That is, increasing the number of inverse temperatures $N_{\beta}$ improves the decoding accuracy, but also increases the computational cost. In contrast, the cost of deterministic algorithms such as the MWPM, CPLEX, and MPS decoders cannot be easily adjusted, and their parallelization is neither straightforward nor efficient.
VII Conclusion
We have demonstrated our decoder for the XZZX code by using simulated annealing (SA). As an initial configuration of SA, we have proposed to employ recovery chains obtained by our decoder which utilizes a kind of greedy matching.
We assumed the code capacity noise, where only the data qubits suffer errors. Under the depolarizing noise and the $Y$-biased noises, we performed the numerical simulation.
We observed that our SA decoder achieves the logical error rate $P_{\mathrm{L}}$, equal to the optimal rate by the CPLEX decoder, which is formulated as an integer programming approach. Moreover, it is more accurate than the MWPM decoder, which cannot take into account the probability of $Y$ errors.
We examined the convergence of the logical error rate for the number of inverse temperatures $N_{\mathrm{\beta}}$ we employ for our SA decoder. In particular, our SA decoder can be interpreted as giving improvement of initial configuration by our greedy matching decoder. Furthermore, we confirmed that the convergence is strongly influenced by methods of constructing a pure error, which are employed to prepare initial configurations. Especially, we proposed the ‘Greedy matching (different)’ method which may produce a different pure error for every invocation, and observed that it leads to markedly faster convergence.
Subsequently, we compared the elapsed times of various decoding algorithms. As a result, our SA decoder is faster than the CPLEX decoder when the physical error rate $p$ is higher. For lower $p$, our SA decoder, if parallelized, is faster than the CPLEX decoder. In summary, we conclude that our parallelizable SA decoder is the fastest among the decoders, including the MPS decoder, which can take into account the probability of $Y$ errors. This parallelization of SAs and its simulation remains a topic for our future work.
We list other future works below. Although our analysis has assumed the code capacity noise model, applying our SA decoder to phenomenological noise and circuit-level noise models is straightforward.
The decoding time may be reduced by improving the MCMC algorithm. Our SA decoder consists of independent SA processes, making it well-suited for parallelized implementation in FPGA (field-programmable gate array). Alternatively, techniques to accelerate the MCMC method could be employed, such as the replica-exchange Monte Carlo method [40] and cluster algorithms including the Wolff algorithm [41]. Furthermore, the Metropolis algorithm can be parallelized by domain (area) decomposition, which is implementable in FPGA [42].
In our SA decoder, we are able to control the trade-off between the decoding accuracy and time, by adjusting the number of Monte Carlo steps, temperature scheduling and convergence criterion, depending on the code distance $d$ and the physical error rate $p$. Finding a suitable set of parameter values may require exhaustive investigation.
The MCMC decoders require only the definitions of stabilizer generators and logical operators of codes, without relying on a decoding graph and, hence, they are applicable to many types of codes, including the color code and quantum LDPC codes [43]. To obtain a pure error to construct an initial error configuration, we can use not only our greedy matching decoder but also other decoders, such as the union-find decoder [44]. For faster convergence, we will pursue ‘good’ initial error configuration.
We plan to perform simulations under actual noise biases $p_{x}:p_{y}:p_{z}$ specific to individual quantum devices. Toward practical use in quantum computing, it is necessary to improve the decoding speed, particularly at lower physical error rates $p$. Our results suggest that our SA decoder is a promising candidate for practical quantum computing.
Acknowledgements.
We would like to thank Keisuke Fujii for his helpful discussions. This work is supported by MEXT Quantum Leap Flagship Program (MEXT Q-LEAP) Grant No. JPMXS0118067394 and JPMXS0120319794, JST COI-NEXT Grant No. JPMJPF2014, and JST Moonshot R&D Grant No. JPMJMS2061.
Appendix A Decoder formulated as integer programming problem
In this section, instead of Eq. (20), we formulate decoding problem as finding the recovery configuration $C$, subject to the constraint $\partial{C}=S$:
$\arg\min_{C}\left\{\mathcal{H}(C)\,\middle|\,\partial{C}=S\right\}.$ (33)
The error configuration $C$ can be represented as the set of one bit binary variables $\{x_{i},y_{i},z_{i}\}_{i}$ where $i=1,\dots,N_{\mathrm{q}}$ indexes the qubits. Namely, $x_{i}=1$ indicates that the qubit $i$ suffers $X$ error, while the variables $y_{i}$ and $z_{i}$ similarly represent $Y$ and $Z$ errors, respectively. Using this binary representation, the syndrome consistency condition $\partial{C}=S$ in Eq. (33) is explicitly expressed as a system of equations: For each face $f$,
$\begin{split}\left(z_{\mathrm{left}(f)}\oplus y_{\mathrm{left}(f)}\right)\oplus\left(z_{\mathrm{right}(f)}\oplus y_{\mathrm{right}(f)}\right)&\\ \,\oplus\left(x_{\mathrm{up}(f)}\oplus y_{\mathrm{up}(f)}\right)\oplus\left(x_{\mathrm{down}(f)}\oplus y_{\mathrm{down}(f)}\right)&=s_{f}\end{split}$ (34)
where $s_{f}\in\{0,1\}$ is the syndrome measurement outcome, as defined in Eq. (2), and the operation $\oplus$ denotes addition modulo two. Under the constraint Eq. (34), the goal is to find the solution $\{x_{i},y_{i},z_{i}\}_{i}$ that minimizes the objective function given by Eq. (10). Here, the variables $n_{x}$, $n_{y}$, $n_{z}$ represent the number of $X$-, $Y$-, $Z$-type errors, respectively, and are given by
$n_{x}=\sum_{i=1}^{N_{\mathrm{q}}}x_{i},\quad n_{y}=\sum_{i=1}^{N_{\mathrm{q}}}y_{i},\quad n_{z}=\sum_{i=1}^{N_{\mathrm{q}}}z_{i}.$
This solution yields a recovery configuration $C$.
This problem falls into the category of (mixed-) integer linear programming, known to be NP-hard. The integer programming formulation exhibits exponential computational complexity with respect to both the code distance $d$ and the physical error rate $p$. Thus, for large values of $d$ or $p$, it is prohibitive to find a solution in a realistic time. To solve the integer linear programming problem, we employ the C++ library, CPLEX (version 22.1.0) [45], whence we call it CPLEX decoder. The use of integer programming solvers for decoding was originally proposed and demonstrated for the color code [46].
Since the CPLEX decoder can take the probabilities $p_{x}$, $p_{y}$, $p_{z}$ into account by Eq. (10), it is capable of achieving the lowest logical error rate among the decoders formulated as the energy minimization problems. Thus, we regard the logical error rate obtained by the CPLEX decoder as a reliable benchmark and thereby evaluate the accuracy of our SA decoder in Sec. V.

### v2

@@
 Abstract
-The XZZX code is a variant of the surface code designed to address biased noise in realistic quantum devices.
-For the XZZX code, we propose a decoder based on simulated annealing (SA).
-Our SA decoder can be readily and efficiently parallelized, by virtue of its simple MCMC-based algorithm.
-To prepare an initial configuration of SA, we propose to employ recovery chains obtained by a decoder which utilizes a kind of greedy matching graph algorithm.
-Although $Z$-biased noise is commonly assumed in real quantum devices, we focus on $Y$-biased noise, for which the minimum-weight perfect matching (MWPM) algorithm fails to decode accurately.
-Our numerical simulation for the code capacity noise model, where only data qubits suffer errors, confirmed that our SA decoder is more accurate than the MWPM decoder.
-Furthermore, our SA decoder attained the accuracy equivalent to that of the optimal decoder formulated by integer programming, called CPLEX decoder.
-In our greedy matching decoder, we randomly determine order of matching pairs of incorrect syndromes that have the same distance.
-This randomness brings about a variety of initial configurations of SA, which leads to faster convergence of our SA decoder.
-By comparing decoding times of our SA decoder, the CPLEX decoder, and matrix product state (MPS) decoder, all of which can handle $Y$-biased noise appropriately, we confirmed that our SA decoder is fastest if parallelized.
-This result implies a potential for combining of our greedy matching and SA decoder for practical use in quantum computing.
-XZZX code, simulated annealing, biased noise, code capacity noise, maximum likelihood decoder, Metropolis algorithm
+The XZZX code is a variant of the surface code tailored to address biased noise in realistic quantum devices.
+We propose a simulated annealing (SA) decoder for the XZZX code.
+Our SA decoder is amenable to parallelization because its MCMC updates are simple and local.
+To initialize SA, we use a recovery configuration produced by our greedy matching decoder.
+Although $Z$-biased noise is commonly assumed in realistic quantum devices, we instead focus on $Y$-biased noise, where MWPM becomes suboptimal because it neglects correlations induced by $Y$ errors.
+Our numerical simulations for the code capacity noise model, where only data qubits suffer errors, show that our SA decoder achieves higher accuracy than the MWPM decoder.
+Furthermore, our SA decoder achieves an accuracy comparable to that of the optimal minimum-energy (MAP-configuration) decoder formulated as an integer programming problem, called the CPLEX decoder.
+In our greedy matching decoder, we randomize the tie-breaking among equal-weight pairs.
+This randomness generates a variety of initial configurations for SA, which can lead to faster convergence of our SA decoder.
+By comparing decoding times of our SA decoder, the CPLEX decoder, and the matrix product state (MPS) decoder, all of which can handle $Y$-biased noise appropriately, we estimate that our SA decoder could be competitive in runtime under an idealized assumption of near-perfect parallel efficiency.
+These results suggest that combining SA with our greedy matching initializer is a practical approach toward fault-tolerant quantum computation.
+XZZX code, simulated annealing, biased noise, code capacity noise model, Metropolis algorithm
 I Introduction
 The ultimate goal of quantum computing is to realize fault-tolerant quantum computation (FTQC) with robust error correction mechanisms.
-For superconducting devices, decoherence time (qubit lifetime) is typically on the order of hundreds of microseconds.
-For this reason, the decoding process must be completed within this time frame.
-Practical decoders require a simple algorithm to be readily implemented in hardware.
-The surface code is one of the most prospective topological error correction codes [1, 2].
-It is experimentally feasible, because it requires measuring only nearest-neighbor data qubits [3, 4, 5, 6, 7].
-In physical systems such as superconducting qubits, optics and ion traps, realistic noise is biased, that is, $Z$ (phase-flip) errors occur more frequently than $X$ (bit-flip) errors.
-For such a biased noise, replacing the stabilizer generators of the surface code leads to alternative codes, such as the tailored surface code (which has only $Y$-type stabilizers) [8] and the XZZX code [9].
-The XZZX code exhibits higher thresholds for $X$-,$Y$- or $Z$-biased noises, compared with the ordinary surface code, which has stabilizer generators consisting of either only $X$ or $Z$ operators, and is known as the CSS (Calderbank–Shor–Steane) surface code [10, 11].
-For the XZZX code, there appeared proposals on its rotated version [12, 13], its cluster state [14] and its physical realization in optical and superconducting devices [15] and neutral atom devices [16].
-The minimum-weight perfect matching (MWPM) decoder [1] can be executed in polynomial time of the number of incorrect syndromes and, hence, is well used for the surface code.
-However, it deals with $X$ and $Z$ errors independently, and is thus less accurate under noises that include $Y$ errors.
-On the other hand, MCMC (Markov chain Monte Carlo) decoder and matrix product state (MPS) decoder [17], known as MLD (maximum likelihood decoder), possess high decoding accuracy by taking into account the probability of $Y$ errors.
-For the $Z$-biased and $Y$-biased noises, the MPS decoder has been researched [18].
-Unfortunately, the MPS decoder has a high computational complexity of order $\mathcal{O}(d^{2})$ for the code distance $d$.
-In contrast, the MCMC decoders have a possibility for fast decoding, since the MCMC algorithms use heuristic approach based on pseudo-random numbers.
-One of the earliest studies [19] on MCMC decoders for the CSS surface code employed parallel tempering, while another study [20] utilized the Metropolis method fixed at the Nishimori temperature.
-For initial configuration of MCMC, they employ error configurations generated by acting stabilizers and logical operators randomly to actual error configuration [19, 12] and recovery chain obtained by the MWPM decoder [20].
-Thus, the aim of these previous works is to establish an accurate threshold for the code.
-To efficiently minimize energy in MCMC methods, simulated annealing (SA) [21] is often used in statistical physics.
-By formulating the decoding problem as a random-bond Ising model (RBIM), SA has been applied to decoders for the CSS surface code [22] and the color code [23].
-Another formulation of SA decoder for the CSS surface code has hamiltonian with penalty terms [24], and is demonstrated in hardware-implemented annealing machine [25, 26, 22].
-The XZZX code has been studied by leveraging the MWPM decoder, the MPS decoder [9] and MCMC decoder [12].
-They investigated the decoding accuracy for $Z$-biased noise, which is typical in real quantum devices.
-Assuming knowing the probabilities of $X,Y,Z$ errors, they incorporate this prior information into weight coefficients of the hamiltonian so as to improve decoding accuracy.
-The research [12] assumes the code capacity noise model, in which only data qubits suffer errors independently and identically.
-To explore configurations efficiently, they carried out the MCMC method at lower inverse temperature than the Nishimori temperature, keeping the weight coefficients invariant in the hamiltonian.
-Their numerical experiment suggests that calculating a dominant term (composed of degeneracy and energy) of the partition function is sufficient to determine the correct logical operator class.
-Thus, their algorithm may be useful to collect this term to utilize it as train data in neural network decoder.
-In the present article, we propose a simulated annealing (SA) decoder for the XZZX code.
-The aim of our research is not to determine an accurate threshold, but rather to establish a decoding algorithm that is moderately accurate and yet sufficiently fast for practical use on a quantum computer.
-We assume the code capacity noise.
-Specifically, our interest is the case that each qubit experiences $Y$-biased noise, which cannot be decoded appropriately by the MWPM decoder.
-As initial configuration for SA, we propose to employ a recovery chain obtained from a decoder which utilizes a greedy matching algorithm.
-We implement the whole algorithms in the C++ language from scratch, in order to attain decoding time as short as possible.
-The rest of the present paper is organized as follows.
+In superconducting qubit platforms, coherence time (i.e., qubit lifetime) is typically on the order of hundreds of microseconds.
+The error correction process must be completed within this time frame.
+Practical decoders therefore favor simple algorithms to be readily implemented in hardware.
+The surface code is one of the most promising topological error correction codes [kitaev, surface_code_towards].
+It is experimentally feasible because it requires only nearest-neighbor interactions between data and ancilla qubits [Zhao, Suppressing_experiment, iOlius, google_surface_code, 17qubit_experiment].
+In physical systems such as superconducting qubits, optical systems, and ion traps, realistic noise is often biased, meaning that $Z$ (phase-flip) errors occur more frequently than $X$ (bit-flip) errors.
+For such biased noise, modifying the stabilizer generators of the surface code leads to alternative codes, such as the tailored surface code (which has only $Y$-type stabilizers) [tuckett_tailoring] and the XZZX code [xzzx_nature].
+The XZZX code achieves higher thresholds under $X$-, $Y$-, or $Z$-biased Pauli noise than the conventional surface code.
+The conventional surface code is a CSS (Calderbank–Shor–Steane) code [calderbank_shor, steane] with stabilizer generators consisting solely of $X$-type or $Z$-type operators.
+For the XZZX code, there have been proposals for its rotated version [xzzx_agnostic, rotated_xzzx], a cluster-state (foliated) construction [foliated_xzzx], numerical simulations for optical and superconducting qubits [Darmawan_xzzx] and neutral-atom qubits [neutral_atom].
+The minimum-weight perfect matching (MWPM) decoder [kitaev] can be executed in polynomial time in the number of syndrome defects and, hence, is widely used for the surface code.
+However, in the standard MWPM approach, since it does not capture the correlations induced by $Y$ errors, its performance can degrade when $Y$ errors are non-negligible.
+On the other hand, an MCMC (Markov chain Monte Carlo) decoder and a matrix product state (MPS) decoder [efficient_MLD] exhibit high decoding accuracy by taking into account the probability of $Y$ errors.
+For $Z$- and $Y$-biased noise, the MPS decoder has been studied [ultrahigh].
+However, for a fixed truncation (bond) dimension $\chi$, the MPS decoder scales as $\mathcal{O}(d^{2})$ (up to polynomial factors in $\chi$).
+In contrast, MCMC-based decoders can be fast in practice because each MCMC update is local and simple, and the computation is easily parallelized.
+One of the earliest studies [Wootton] on MCMC decoders for the CSS surface code employed parallel tempering, while another study [Hutter] utilized the Metropolis method fixed at the Nishimori inverse temperature.
+For an initial error configuration of MCMC, they employed error configurations generated by acting stabilizers and logical operators randomly on the true error configuration [Wootton, xzzx_agnostic] and a recovery chain obtained by the MWPM decoder [Hutter].
+Thus, the aim of these previous works was to determine the threshold of the code accurately.
+To efficiently minimize energy in MCMC methods, simulated annealing (SA) [Kirkpatrick] is often used in statistical physics.
+By formulating the decoding problem as a random-bond Ising model (RBIM), SA has been applied to decoders for the CSS surface code [takeuchi_comparative_decoder] and the color code [takada].
+Another formulation of an SA decoder for the CSS surface code has a Hamiltonian with penalty terms [fujii_penalty], and was demonstrated in a hardware-implemented annealing machine [fujisaki_DA, fujisaki_circuit-level, takeuchi_comparative_decoder].
+The XZZX code has been studied using the MWPM decoder, the MPS decoder [xzzx_nature], and the MCMC-based decoders [xzzx_agnostic].
+They investigated the decoding accuracy for $Z$-biased noise, which is typical in quantum hardware platforms.
+Assuming knowledge of the probabilities of $X,Y,Z$ errors, the authors incorporated this prior information into the weight coefficients of the Hamiltonian so as to improve decoding accuracy.
+Ref.
+[xzzx_agnostic] assumes the code capacity noise model, in which only data qubits suffer independently and identically distributed errors.
+To explore configurations efficiently, they carried out the MCMC method at a lower inverse temperature than the Nishimori inverse temperature while keeping the weight coefficients invariant in the Hamiltonian.
+Their numerical results suggest that extracting a dominant term of the partition function is sufficient to determine the correct logical operator class.
+Thus, their algorithm may be useful for collecting this term, which could then be used as training data for a neural-network decoder.
+In this paper, we propose a simulated annealing (SA) decoder for the XZZX code.
+Our goal is not to determine the threshold as accurately as possible, but rather to develop a decoding algorithm that is reasonably accurate, yet sufficiently fast for practical use on a quantum computer.
+We assume the code capacity noise model.
+Specifically, we are interested in the case where each qubit experiences $Y$-biased noise, a regime in which the MWPM decoder becomes significantly suboptimal because it neglects correlations induced by $Y$ errors.
+To initialize SA, we employ a recovery configuration produced by a greedy matching decoder.
+All algorithms are implemented in C++ to minimize decoding time.
+The rest of this paper is organized as follows.
 In Sec.
 II, we review the XZZX code.
-For the XZZX code, we describe principle of error detection and correction in Sec.
-III and decoding algorithms, including our SA decoders in Sec.
+For the XZZX code, we describe the principles of error detection and correction in Sec.
+III, and then present decoding algorithms which include our SA decoders in Sec.
 IV.
-We show numerical result and argue decoding accuracy and time in Sec.
+We show numerical results in Sec.
 V and discuss the benefits of our SA decoder in Sec.
 VI.
-In Sec.
-VII we conclude.
+We conclude in Sec.
+VII.
 II The XZZX code
-In this section, we briefly review the XZZX surface code [9]. We consider a rotated square lattice with open boundaries [27]. In the lattice, data qubits reside on the vertices.
+In this section, we briefly review the XZZX surface code [xzzx_nature]. We consider a rotated square lattice with open boundaries [kitaev-open_boundary]. Data qubits reside on the vertices.
 (a) Refer to caption
 (b) Refer to caption(c) Refer to caption
-Figure 1: The XZZX code for the code distance $d=5$.
-(a) Stabilizer generators $G_{f}$, showing four-body in the bulk and three-body operators at the boundaries.
-(b) Logical operator $L_{Z}$ (c) Logical operator $L_{X}.$
+Figure 1: The XZZX code for the code distance $d=5$: (a) Stabilizer generators $G_{f}$, showing four-body operators in the bulk and three-body operators at the boundaries.
+(b) Logical operator $L_{Z}$.
+(c) Logical operator $L_{X}$.
 We illustrate stabilizer generators for the XZZX surface code in Fig. 1(1). The stabilizer generator acting around a face $f$ is defined as
 $G_{f}\coloneqq X_{\mathrm{left}(f)}\,Z_{\mathrm{up}(f)}\,Z_{\mathrm{down}(f)}\,X_{\mathrm{right}(f)}$ (1)
-whose form is the origin of the name ‘XZZX code’.
-This stabilizer generator is a four-body operator in bulk and a three-body operator at the boundaries.
+whose form is the origin of the name “XZZX code”.
+This stabilizer generator is a four-body operator in the bulk and a three-body operator at the boundaries.
 The XZZX code has only one type of stabilizer generators, unlike the CSS surface code which has two types of generators.
-Since each of the stabilizer generator includes both $X$ and $Z$ operators as shown in Eq.
-(1), the XZZX code is not CSS-type.
+Since each stabilizer generator includes both $X$ and $Z$ operators as shown in Eq.
+(1), the XZZX code is not a CSS code.
+At the center of every face $f$, an ancilla qubit is placed to measure the stabilizer generator $G_{f}$ and detect errors.
+We define logical operators $L_{Z}$ ($L_{X}$) as a Pauli chain connecting the left and right (top and bottom) boundaries, respectively.
+Fig.
+1(1) and (1) depict examples of the logical operators $L_{Z}$ and $L_{X}$, respectively.
+We further define $L_{Y}\coloneqq\mathrm{i}\,L_{Z}\,L_{X}$.
+Throughout this paper, we ignore global phases of Pauli operators (i.e., overall factors $\pm 1$ and $\pm\mathrm{i}$ are identified).
+Let $\overline{\mathcal{P}}$ denote the $N_{\mathrm{q}}$-qubit Pauli group modulo global phases.
+In particular, we identify $L_{Y}$ with $L_{Z}L_{X}$ in $\overline{\mathcal{P}}$.
+We identify Pauli chains that differ by multiplication of a stabilizer operator in the stabilizer group $\mathcal{G}\coloneqq\langle G_{f}\mid f:\text{face}\rangle$.
+More precisely, for $P\in\{I,X,Y,Z\}$, we define the logical Pauli operator as the equivalence class $L_{P}]$ modulo stabilizers, i.e., $L_{P}\sim GL_{P}$ for any $G\in\mathcal{G}$.
+The code distance $d$ is defined as the minimum weight of a nontrivial logical operator (equivalently, the length of the shortest Pauli chain representing $L_{Z}$ or $L_{X}$).
+For the surface codes including the XZZX code, this distance coincides with the linear size of the lattice.
+The number of data qubits is $N_{\mathrm{q}}=d^{2}+(d-1)^{2}$.
 The number of stabilizer generators is equal to the number of faces, i.e., $N_{\mathrm{stabilizer}}=2d(d-1)=\mathcal{O}(d^{2})$.
-At the center of every face $f$, an ancilla qubit is located to detect errors. For each ancilla qubit, we measure the stabilizer generator $G_{f}$.
-We define logical operators $L_{Z}$ ($L_{X}$) as a chain connecting the left and right (upper and lower) boundaries, respectively.
-Fig.
-1(1),(1) depict examples of the logical operators $L_{Z}$ and $L_{X}$.
-Furthermore, we define logical operator $L_{Y}\coloneqq\mathrm{i}\,L_{Z}\,L_{X}$.
-As long as the opposite boundaries are connected, we regard these chains as logical operators, namely, allow the freedom of a stabilizer operator in the stabilizer group $\mathcal{G}\coloneqq\langle G_{f}\,|\,f:\text{face}\rangle$.
-More precisely, for $P\in\{I,X,Z,Y\}$, the logical operator is defined as equivalence class $L_{P}]$ by imposing invariance under the action of the stabilizer group $\mathcal{G}$ in the sense that $G\,L_{P}]=[L_{P}]$ for any $G\in\mathcal{G}$.
-The code distance $d$ is defined as the length of the shortest representation of the logical operators $L_{Z}$ and $L_{X}$.
-For the surface codes including the XZZX code, it is the linear length of the lattice.
-The number of data qubits is $N_{\mathrm{q}}=d^{2}+(d-1)^{2}$.
-The XZZX code is advantageous under biased noises.
-That is, the XZZX code has threshold $p_{\mathrm{th}}=0.5$ for pure $X$, $Y$, $Z$ noises, because it is equivalent to classical repetition codes; On the other hand, the CSS surface code has $p_{\mathrm{th}}=0.109$ for the pure $X,Z$ noises [1] and $p_{\mathrm{th}}=0.5$ for the pure $Y$ noise [8].
-III Principle of Error detection and correction
-Throughout the present paper, error configuration means the information that each data qubit suffers which $I$ (no error), $X$, $Z$ or $Y$ error.
+The XZZX code is advantageous under biased noise.
+That is, the XZZX code has a threshold $p_{\mathrm{th}}=0.5$ for pure $X$, $Y$, or $Z$ noise in the code capacity noise model, reflecting its mapping to a classical repetition code [xzzx_nature]; on the other hand, the CSS surface code has $p_{\mathrm{th}}=0.109$ for the pure $X$ or $Z$ noise [kitaev] and $p_{\mathrm{th}}=0.5$ for the pure $Y$ noise [tuckett_tailoring].
+III Principles of error detection and correction
+Throughout this paper, an error configuration $C\in\overline{\mathcal{P}}$ specifies, for each data qubit, whether it experiences no error ($I$) or a Pauli error $X$, $Y$, or $Z$.
 Fig.
@@
 The set of vertices $\mathcal{V}_{\bullet}$
-The vertices are the centers of faces with the color $\bullet$ and are the ancillas for the stabilizer measurement.
+The vertices are the centers of faces with the color $\bullet$; each vertex corresponds to an ancilla qubit used to measure the stabilizer.
 The set of edges $\mathcal{E}_{\bullet}$
-The edges which connect nearby vertices in $\mathcal{V}_{\bullet}$. The edge centers represent errors which suffer data qubits.
-The set of edges which suffer errors is called error chain.
-There are two types of error chains $C_{\mathrm{white}}\subset\mathcal{E}_{\mathrm{white}}$ and $C_{\mathrm{gray}}\subset\mathcal{E}_{\mathrm{gray}}$, which connect centers of white and gray face, respectively, drawn with the pink and purple lines in Fig.
+The edges connect nearest-neighbor vertices in $\mathcal{V}_{\bullet}$.
+Each edge is associated with the data qubit located at its midpoint; an error on that data qubit is represented on the corresponding edge.
+The set of edges on which errors occur is called an error chain.
+There are two types of error chains $C_{\mathrm{white}}\subset\mathcal{E}_{\mathrm{white}}$ and $C_{\mathrm{gray}}\subset\mathcal{E}_{\mathrm{gray}}$, which connect centers of white and gray faces, respectively, and are drawn with the pink and purple lines in Fig.
 2.
-For both the error chains $C_{\mathrm{white}}$ and $C_{\mathrm{gray}}$, the horizontal (vertical) edges represent $Z$ ($X$) errors, respectively.
-Solely on the intersection points of the horizontal and vertical edges in the different types of error chains, $Y$ errors appear.
-Figure 2: An example of the error configuration.
-The circles indicate qubits which suffer errors with error type $X$,$Y$,$Z$.
-(The qubits with no errors are not drawn.) The error configuration is expressed as two types of error chains $C_{\mathrm{white}}$ and $C_{\mathrm{gray}}$, connecting the white faces and gray faces and drawn with the pink and purple lines, respectively.
-The endpoints of these error chains are captured by the incorrect syndromes, drawn with the green squares.
-In both the error chains $C_{\mathrm{white}}$ and $C_{\mathrm{gray}}$, the $Z$ ($X$) errors reside on horizontal (vertical) edges, respectively.
-The $Y$ error appears solely on the intersection point of both the error chains $C_{\mathrm{white}}$ and $C_{\mathrm{gray}}$.
-For each face $f$, we measure the stabilizer generator $G_{f}$ and thereby obtain a measurement outcome $s_{f}\in\{0,1\}$. For an error chain $C_{\bullet}$, it holds the eigenvalue equation
-[…]
-[…]
+In this paper, to avoid confusion, $C$ (without a subscript) denotes an error configuration, whereas $C_{\bullet}$ (with $\bullet\in\{\mathrm{white},\mathrm{gray}\}$) denotes an error chain.
+For both the error chains $C_{\mathrm{white}}$ and $C_{\mathrm{gray}}$, the horizontal and vertical edges represent the presence of $Z$- and $X$-type Pauli components, respectively (i.e., $Z$ or $Y$ on a horizontal edge, and $X$ or $Y$ on a vertical edge).
+A $Y$ error occurs exactly at the intersection point of both the error chains $C_{\mathrm{white}}$ and $C_{\mathrm{gray}}$.
+With a slight abuse of notation, we use the same symbol $C_{\bullet}$ to denote the corresponding Pauli operator.
+Figure 2: An example of the error configuration $C$.
+Circles indicate data qubits with $X$, $Y$, or $Z$ errors; qubits without errors are omitted.
+The error configuration can be represented as two types of error chains $C_{\mathrm{white}}$ and $C_{\mathrm{gray}}$, connecting the white faces and gray faces and drawn with the pink and purple lines on the corresponding decoding graphs $(\mathcal{V}_{\mathrm{white}},\mathcal{E}_{\mathrm{white}})$ and $(\mathcal{V}_{\mathrm{gray}},\mathcal{E}_{\mathrm{gray}})$, respectively.
+The endpoints of these error chains are captured by the syndrome defects, drawn with the green squares.
+In both the error chains $C_{\mathrm{white}}$ and $C_{\mathrm{gray}}$, the $Z$- and $X$-type Pauli components correspond to operators on data qubits located at the centers of horizontal and vertical edges, respectively.
+A $Y$ error occurs precisely on a data qubit where both the $Z$- and $X$-type components are present (i.e., where the corresponding horizontal and vertical edges are both occupied).
+For each face $f$, we measure the stabilizer generator $G_{f}$ and thereby obtain a measurement outcome $s_{f}\in\{0,1\}$.
+For an error chain (viewed as a Pauli operator) $C_{\bullet}$, the outcome is determined by whether $C_{\bullet}$ commutes or anti-commutes with $G_{f}$:
+$G_{f}\,C_{\bullet}=(-1)^{s_{f}}\,C_{\bullet}\,G_{f},$ (2)
+i.e., $s_{f}=0$ if $G_{f},C_{\bullet}]=0$ and $s_{f}=1$ if $\{G_{f},C_{\bullet}\}=0$. We refer to a face $f$ with $s_{f}=1$ as a syndrome defect.
+A syndrome defect appears precisely at an endpoint of the error chain $C_{\bullet}$.
+More precisely, $s_{f}=1$ if and only if $v_{f}\in\partial C_{\bullet}$, where $v_{f}$ denotes the vertex (i.e., face center) corresponding to the face $f$, and $\partial C_{\bullet}$ denotes the set of vertices in $\mathcal{V}_{\bullet}$ that have odd incidence with the edge set $C_{\bullet}$ (equivalently, the endpoints of the chain $C_{\bullet}$).
 In Fig.
-2, incorrect syndromes are drawn with green squares.
-The incorrect syndromes capture the endpoints of $C_{\mathrm{white}}$ and $C_{\mathrm{gray}}$.
-In the endpoint of horizontal (vertical) edges, $Z$ ($X$) errors anti-commute with $X$ ($Z$) operators acting on the left and right (up and down) qubits in the stabilizer generator $G_{f}$ and hence can be detected.
-For an error chain $C_{\bullet}$, the set of face centers whose ancillas possess incorrect syndromes is called syndrome and denoted by
-$S_{\bullet}\coloneqq\left\{(\text{center of face $f$})\in\mathcal{V}_{\bullet}\,|\,s_{f}=1\right\}.$ (3)
-The syndrome consistency can then be expressed as
+2, syndrome defects are drawn with green squares.
+They capture the endpoints of $C_{\mathrm{white}}$ and $C_{\mathrm{gray}}$.
+At such endpoints of horizontal and vertical edges, $Z$- and $X$-type error components anti-commute with $X$ and $Z$ operators acting on the left/right and up/down qubits in the stabilizer generator $G_{f}$, respectively, and hence can be detected.
+For an error chain $C_{\bullet}$, the set of face centers whose ancilla qubits possess syndrome defects is called the syndrome and denoted by
+$S_{\bullet}\coloneqq\left\{v_{f}\in\mathcal{V}_{\bullet}\,\middle|\,s_{f}=1\right\}.$ (3)
+The syndrome-consistency condition can then be expressed as
 $\partial{C}_{\mathrm{white}}=S_{\mathrm{white}},\qquad\partial{C}_{\mathrm{gray}}=S_{\mathrm{gray}}.$ (4)
-The two error chains $C_{\mathrm{white}}$ and $C_{\mathrm{gray}}$ contain the same information as the error configuration CC [^1] In the present article, to distinguish an error configuration $C$, we denote an error chain by $C_{\bullet}$ with $\bullet\in\{\mathrm{white},\mathrm{gray}\}$.
-.
+The two error chains $C_{\mathrm{white}}$ and $C_{\mathrm{gray}}$ contain the same information as the error configuration $C$.
 We express the correspondence as $C=(C_{\mathrm{white}},C_{\mathrm{gray}})$ and then define the tuple $\partial{C}\coloneqq(\partial{C_{\mathrm{white}}},\partial{C_{\mathrm{gray}}})$.
 Similarly, we define the tuple $S\coloneqq(S_{\mathrm{white}},S_{\mathrm{gray}})$.
-Using this abbreviated notation, the syndrome consistency of Eq.
+Using this abbreviated notation, the syndrome-consistency condition of Eq.
 (4) reads $\partial{C}=S$.
-The actual error configuration $C$ is unknown; The only information available is the syndrome $S$, from which decoders infer a recovery configuration $C$ such that $\partial{C}=S$.
-Decoders can be classified into two types: those that identify the maximum likelihood error (MLE) configuration and those that identify the maximum-likelihood logical operator class.
-The latter is referred to as the maximum likelihood decoder (MLD).
-The examples of the MLE decoder are the MWPM decoder (Sec.
-IV.2) and our greedy matching decoder (Sec.
-IV.3).
-The examples of MLD are our SA decoders (Sec.
-IV.1, Sec.
-IV.4), decoder formulated as integer programming (Sec.
-A) and the MPS decoder [17].
-In general, the second type is more accurate than the first type, because it can take into account the degeneracy of error configurations with the same energy.
+The true error configuration $\bar{C}$ is unknown; the only available information is the syndrome $S$, from which decoders infer a recovery configuration $C$ such that $\partial{C}=S$.
 IV Decoding algorithms
@@
 As a simplified version, in Sec.
-IV.3, we propose a greedy matching decoder, exploited to construct an initial configuration of the SA decoder.
+IV.3, we propose a greedy matching decoder, exploited to construct an initial configuration for the SA decoder.
 IV.1 Simulated annealing (SA) decoder
-We assume the code capacity noise where only the data qubits suffer errors.
-Specifically, each data qubit suffers an error of $X$-, $Y$-, or $Z$-type, independently and identically obeying the probabilities $p_{x}$, $p_{y}$, $p_{z}$, respectively.
-Any type of error occurs with the probability $p\coloneqq p_{x}+p_{y}+p_{z}$, called physical error probability.
-This noise for each data qubit $\rho$ is expressed as quantum channel
+We assume the code capacity noise model where only data qubits are noisy and syndrome measurements are perfect.
+Specifically, each data qubit suffers an error of $X$-, $Y$-, or $Z$-type, independently and identically distributed with the probabilities $p_{x}$, $p_{y}$, $p_{z}$, respectively.
+A non-identity Pauli error occurs with the probability $p\coloneqq p_{x}+p_{y}+p_{z}$, called the physical error rate.
+The noise acting on each data qubit state $\rho$ is given by the channel
 $\mathcal{E}(\rho)=(1-p)\,\rho\,+\,p_{x}X\rho X+p_{y}Y\rho Y+p_{z}Z\rho Z.$ (5)
-For an error configuration $C$, let $n_{x},n_{y},n_{z}$ denote the numbers of qubits which suffer $X,Y,Z$ errors, respectively. The probability of $C$ is then expressed as
-$\operatorname{Prob}(C)$ $\coloneqq$ ${p_{x}}^{n_{x}}\,{p_{y}}^{n_{y}}\,{p_{z}}^{n_{z}}\,(1-p)^{N_{\mathrm{q}}-(n_{x}+n_{y}+n_{z})}$ (6)
+For an error configuration $C$, let $n_{x},n_{y},n_{z}$ denote the numbers of qubits which suffer $X,Y,Z$ errors, respectively. The probability of $C$ is then given by
+$\Pr(C)$ $\coloneqq$ ${p_{x}}^{n_{x}}\,{p_{y}}^{n_{y}}\,{p_{z}}^{n_{z}}\,(1-p)^{N_{\mathrm{q}}-(n_{x}+n_{y}+n_{z})}$ (6)
 $=$ $(1-p)^{N_{\mathrm{q}}}\,\left(\frac{p_{x}}{1-p}\right)^{n_{x}}\,\left(\frac{p_{y}}{1-p}\right)^{n_{y}}\,\left(\frac{p_{z}}{1-p}\right)^{n_{z}}$
@@
 $\pi(C)$ $\coloneqq$ $\left(\frac{p_{x}}{1-p}\right)^{n_{x}}\,\left(\frac{p_{y}}{1-p}\right)^{n_{y}}\,\left(\frac{p_{z}}{1-p}\right)^{n_{z}}$ (7)
-$=$ $\left(\frac{p}{1-p}\right)^{\alpha_{x}n_{x}+\alpha_{y}n_{y}+\alpha_{z}n_{z}}$ (8)
-is a factor relevant to $p_{x},p_{y},p_{z}$. From Eqs. (7),(8), the coefficients $\alpha_{x},\alpha_{y},\alpha_{z}$ are expressed as
-$\alpha_{\mu}=\frac{\log{\frac{p_{\mu}}{1-p}}}{\log{\frac{p}{1-p}}}$ (9)
-where $\mu\in\{x,y,z\}$. Assuming that we know or estimate $p_{x}$, $p_{y}$, $p_{z}$ enables optimal decoding, as discussed below.
-We define energy of the error configuration $C$ as the weighted sum of the numbers [^2] The numbers $n_{x}$, $n_{y}$, $n_{z}$ are calculated as follows.
+is the (unnormalized) Boltzmann weight of configuration $C$, i.e., $\Pr(C)$ up to the constant factor $(1-p)^{N_{\mathrm{q}}}$.
+For convenience, we re-express Eq. (7) using a common base p/(1p)p/(1-p):
+$\pi(C)$ $=$ $\left(\frac{p}{1-p}\right)^{\alpha_{x}n_{x}+\alpha_{y}n_{y}+\alpha_{z}n_{z}}$ (8)
+with
+$\alpha_{\mu}\coloneqq\frac{\log\!{\left(\frac{p_{\mu}}{1-p}\right)}}{\log\!{\left(\frac{p}{1-p}\right)}}$ (9)
+where $\mu\in\{x,y,z\}$.
+We assume $0<p<1/2$, so that $\log(p/(1-p))<0$.
+Knowing or accurately estimating $p_{x}$, $p_{y}$, $p_{z}$ enables optimal minimum-energy (MAP-configuration) decoding under the dominant-term approximation discussed below.
+We define the energy of the error configuration $C$ as the weighted sum of the numbers [^1] The numbers $n_{x}$, $n_{y}$, $n_{z}$ are calculated as follows.
 We formulate and implement the Metropolis method with bit representation taking values in the set $\{0,1\}$, instead of Ising spins taking values in the set $\{1,-1\}$.
-In an error configuration, an error that suffers each data qubit is represented by two bits.
-In the two bits, the upper and lower bit means for $X$ and $Z$ error, respectively.
+In an error configuration, the error on each data qubit is represented by two bits.
+The upper and lower bits represent an $X$ and a $Z$ error, respectively.
 Thus, $I$, $Z$, $X$, $Y$ errors are represented as $00$, $01$, $10$, $11$, respectively.
@@
 Using Eq. (10), Eq. (8) is expressed as
-$\pi(C)=\left(\frac{p}{1-p}\right)^{\mathcal{H}(C)}=e^{-\beta\,\mathcal{H}(C)}$ (11)
-where $\beta$ is inverse temperature, defined as [^3] The Nishimori temperature includes the coefficient $1/2$ in the formulation by Ising spins [28, 29].
-Instead, we employ the formulation by bit representation, and hence, this coefficient does not appear.
-$\beta\coloneqq-\log\left(\frac{p}{1-p}\right).$ (12)
-Let us define an equivalence class for an error configuration $C$ via the equivalence class $\cdot\,$].
-Given a syndrome $S$, we can construct a pure error, i.e., an error configuration $T(S)$ consistent with the syndrome $S$.
-Using any pure error $T(S)$, any error configuration $C$ consistent with the syndrome $S$ can be decomposed as
-$C=T(S)\,L_{P}\,G$ (13)
-where $P\in\{I,X,Y,Z\}$ and a stabilizer operator $G\in\mathcal{G}$ and vice versa [30].
-Left-multiplying both sides of Eq. (13) by $T(S)$ yields
-$T(S)\,C]=[L_{P}].$ (14)
-Then, the set of error configurations $C$ satisfying Eq. (14) is defined as
-$\operatorname{Rep}_{T(S)}(L_{P})$ $\coloneqq$ $\left\{C\,\middle|\,[T(S)\,C]=[L_{P}]\right\}$ (15)
-$=$ $\left\{T(S)\,L_{P}\,G\,\middle|\,G\in\mathcal{G}\right\},$ (16)
-which is the set of representatives of another equivalence class, referred to as class for the logical operator $L_{P}$ based on $T(S)$.
-We employ a pure error $T(S)$ as output of some decoder.
-Specifically, we utilize our greedy matching decoder, elaborated in Sec.
-IV.3.
-The probability of the class for the logical operator $L_{P}$ based on $T(S)$ is given by
-$\mathrm{P}_{L_{P}}\coloneqq\max_{C}\left\{\operatorname{Prob}(C)\,\middle|\,C\in\operatorname{Rep}_{T(S)}(L_{P})\right\}.$ (17)
-Eq. (17) is expressed by using Eqs. (6),(11) as
-$\mathrm{P}_{L_{P}}\propto e^{-\beta\,E_{L_{P}}}$ (18)
+$\pi(C)=\left(\frac{p}{1-p}\right)^{\mathcal{H}(C)}=e^{-\beta_{\mathrm{N}}\,\mathcal{H}(C)},$ (11)
 where
-$E_{L_{P}}$ $\coloneqq$ $\min_{C}\left\{\mathcal{H}(C)\,\middle|\,C\in\operatorname{Rep}_{T(S)}(L_{P})\right\}$ (19)
-$=$ $\min_{G\in\mathcal{G}}\mathcal{H}\left(T(S)\,L_{P}\,G\right)$
+$\beta_{\mathrm{N}}\coloneqq-\log\!\left(\frac{p}{1-p}\right)\qquad\text{for }0<p<1/2.$ (12)
+We refer to $\beta_{\mathrm{N}}$ as the Nishimori inverse temperature corresponding to the physical error rate $p$.
+Conditioned on the syndrome-consistency $\partial C=S$, we consider the posterior $\Pr(C\mid S)\propto\pi(C)$, which follows from Eq.
+(6).
+Given a syndrome $S$, let
+$\Omega_{S}\coloneqq\left\{C\in\overline{\mathcal{P}}\,\middle|\,\partial C=S\right\}$ (13)
+denote the set of syndrome-consistent configurations.
+Fix $R(S)\in\Omega_{S}$, referred to as a reference configuration [^2] Such a fixed $R(S)\in\Omega_{S}$ is often called a pure error in the literature [fujii_book]..
+Then, any error configuration $C\in\Omega_{S}$ can be decomposed as
+$C=R(S)\,L_{P}\,G$ (14)
+for some $P\in\{I,X,Y,Z\}$ and some $G\in\mathcal{G}$ [Poulin].
+Left-multiplying both sides of Eq. (14) by $R(S)$ and using $R(S)^{2}=I$ (up to phase) gives
+$R(S)\,C=L_{P}\,G.$ (15)
+The product R(S)CR(S)\,C has a trivial syndrome (i.e., $\partial(R(S)\,C)=\emptyset$), so its equivalence class [R(S)C][R(S)\,C] is well-defined (as in Sec.
+II) and equals one of $L_{P}]$ with $P\in\{I,X,Y,Z\}$.
+We then define, for each $P\in\{I,X,Y,Z\}$, a coset
+$\operatorname{\mathcal{C}}_{R(S)}(L_{P})\coloneqq\left\{C\in\Omega_{S}\,\middle|\,[R(S)\,C]=[L_{P}]\right\}.$ (16)
+Note that $L_{P}]$ is an equivalence class in $\overline{\mathcal{P}}/\mathcal{G}$, whereas $\operatorname{\mathcal{C}}_{R(S)}(L_{P})$ is a set of concrete configurations $C\in\Omega_{S}$; these objects live in different spaces.
+When $S=\emptyset$ and we choose $R(S)=I^{\otimes N_{\mathrm{q}}}$, for each $P\in\{I,X,Y,Z\}$, $\operatorname{\mathcal{C}}_{R(S)}(L_{P})$ corresponds to the logical operator class $L_{P}]$.
+The four cosets $\{\operatorname{\mathcal{C}}_{R(S)}(L_{P})\}_{P\in\{I,X,Y,Z\}}$ form a partition of $\Omega_{S}$:
+$\Omega_{S}=\operatorname{\mathcal{C}}_{R(S)}(L_{I})\,\dot{\cup}\,\operatorname{\mathcal{C}}_{R(S)}(L_{X})\,\dot{\cup}\,\operatorname{\mathcal{C}}_{R(S)}(L_{Y})\,\dot{\cup}\,\operatorname{\mathcal{C}}_{R(S)}(L_{Z}),$
+[…]
+$\operatorname{\mathcal{C}}_{R(S)}(L_{P})=\left\{R(S)\,L_{P}\,G\,\middle|\,G\in\mathcal{G}\right\}.$ (17)
+We define the maximum (unnormalized) Boltzmann weight within each coset as
+$\pi^{\max}_{L_{P}}\coloneqq\max_{C\in\operatorname{\mathcal{C}}_{R(S)}(L_{P})}\pi(C).$ (18)
+A configuration $C$ that attains the maximum in Eq.
+(18) is the most likely configuration in the coset $\operatorname{\mathcal{C}}_{R(S)}(L_{P})$ under the posterior conditioned on $\partial C=S$.
+Using Eq.
+(11), Eq.
+(18) is expressed as
+$\pi^{\max}_{L_{P}}=e^{-\beta_{\mathrm{N}}\,E^{\min}_{L_{P}}}$ (19)
+where
+$E^{\min}_{L_{P}}$ $\coloneqq$ $\min_{C\in\operatorname{\mathcal{C}}_{R(S)}(L_{P})}\mathcal{H}(C)$ (20)
+$=$ $\min_{G\in\mathcal{G}}\mathcal{H}\left(R(S)\,L_{P}\,G\right)$
 which follows from Eq. (17).
-We determine the most likely class for the logical operator $L_{P^{*}}$ which maximizes the error probabilities $\mathrm{P}_{L_{P}}$:
-$P^{*}\coloneqq\operatorname*{\arg\max}_{P\in\{I,X,Y,Z\}}\mathrm{P}_{L_{P}}=\operatorname*{\arg\min}_{P\in\{I,X,Y,Z\}}E_{L_{P}},$ (20)
-which follows from Eq.
+We determine $P^{\ast}$ by maximizing $\pi^{\max}_{L_{P}}$ (equivalently, minimizing $E^{\min}_{L_{P}}$):
+$P^{\ast}\coloneqq\operatorname*{\arg\max}_{P\in\{I,X,Y,Z\}}\pi^{\max}_{L_{P}}=\operatorname*{\arg\min}_{P\in\{I,X,Y,Z\}}E^{\min}_{L_{P}},$ (21)
+which follows from Eq. (19). This decoding formulation is based on domain-wall constructions in spin-glass models [fujii_book, §3.5] [hukushima_domain_wall, Chubb_stat_model].
+Ideally, one would perform maximum-likelihood (ML) decoding over logical operator cosets by comparing the coset weights.
+We instead use the maximum a posteriori (MAP): we approximate each coset weight $\sum_{C\in\operatorname{\mathcal{C}}_{R(S)}(L_{P})}\pi(C)$ by its largest term $\pi^{\max}_{L_{P}}$ defined in Eq.
 (19).
-This decoding principle is based on the theory of the domain-wall of spin-glass models [31, §3.5] [32, 33].
-For each $P\in\{I,X,Y,Z\}$, to calculate $E_{L_{P}}$ in Eq.
-(19), we employ simulated annealing (SA) [21].
+This reduces ML decoding over cosets to a minimum-energy search within each coset.
+For each $P\in\{I,X,Y,Z\}$, to estimate $E^{\min}_{L_{P}}$ in Eq.
+(20), we employ simulated annealing (SA) [Kirkpatrick].
 For our SA, in the decomposition of Eq.
-(13), $C_{L_{P}}\coloneqq T(S)\,L_{P}$ is set as an initial configuration of SA, and the stabilizer operator $G$ is updated by the Metropolis method to explore the ground state energy $E_{L_{P}}$.
-Although it is rigorously proved that SA converges to the optimal solution regardless of the initial configuration if we take an infinite time [34], we expect that in realistic finite time, the initial configuration affects the convergence speed of SA.
+(14), $C_{L_{P}}\coloneqq R(S)\,L_{P}$ is set as an initial configuration for SA, and the stabilizer component $G$ is updated by the Metropolis method to search for the ground state energy $E^{\min}_{L_{P}}$.
+In most of our simulations, we take $R(S)$ to be the recovery configuration returned by our greedy matching decoder (Sec.
+IV.3).
+Although it has been rigorously proved that SA converges to the optimal solution (under suitable cooling schedules) in the infinite-time limit [Geman_Geman], regardless of the initial configuration, in practical finite time, the initial configuration can significantly affect the convergence speed of SA.
 We examine this effect in Sec.
@@
 We give our SA decoder in Algorithm 1.
-To acquire the lowest possible energy $E_{L_{P}}$, we carry out simulated annealing multiple times (say $N_{\mathrm{SA}}$ times) for the identical error configuration $C_{L_{P}}$ in Lines 78 of Algorithm 1, and select the minimum energy among them in Line 9.
-In Line 14, we finally obtain the recovery configuration $C_{L_{P^{*}}}=T\,L_{P^{*}}$.
-[^4] As a recovery error configuration, any configuration $C\in\operatorname{Rep}_{T(S)}(L_{P^{*}})$ can be selected.
-It should be noted that, in the nested for loops in Lines 5 and 7 in Algorithm 1, each iteration is an independent SA task and, hence, these tasks are obviously parallelizable.
-The total number of iterations in these loops are $4$ and $N_{\mathrm{SA}}$ and, therefore, if parallelized, the elapsed time in parallel execution is expected to be the elapsed time in sequential execution divided by $4N_{\mathrm{SA}}$.
-This parallelized implementation is left to our future work.
+To estimate $E^{\min}_{L_{P}}$ effectively, we run simulated annealing multiple times (say, $N_{\mathrm{SA}}$ times) starting from the identical error configuration $C_{L_{P}}$ (Lines 78 of Algorithm 1), and select the minimum energy across the runs (Line 9).
+In Line 14, we output the recovery configuration $C_{L_{P^{\ast}}}=R\,L_{P^{\ast}}$.
+[^3] We use SA only to estimate $E^{\min}_{L_{P}}$; the output recovery configuration is chosen as $R\,L_{P^{\ast}}$, since any configuration of the inferred logical coset $\operatorname{\mathcal{C}}_{R(S)}(L_{P})$ is equivalent up to stabilizers.
+Note that, in the nested for loops in Lines 5 and 7 in Algorithm 1, each iteration is an independent SA task and, hence, these tasks are trivially parallelizable.
+Implementing this parallelization (and quantifying the overhead) is left for future work.
 Algorithm 1 invokes the simulated annealing, shown in Algorithm 2.
+Note that the simulated_annealing function is called with a copy of the initial configuration.
 Algorithm 2 performs the Metropolis method, given in Algorithm 3.
-Line 9 in Algorithm 3 flips spins, corresponding to errors on qubits, by acting a stabilizer generator [^5] These spins are implemented by bits.
-Hence, the multiplication of error configurations $C\,G_{f}$ is implemented as addition modulo two.
-in order to keep satisfying the syndrome consistency condition $\partial{C}=S$ throughout the entire SA algorithm.
-For the physical error probability $p$, Eq.
-(12) gives the Nishimori temperature $\beta_{\mathrm{nishimori}}$ where actual error configuration and recovery configuration equally fluctuate, and we can incorporate the degeneracy of error configurations with the same energy.
-Hence, it is optimal for decoding as a fixed temperature [28, 29, 33].
-In Algorithm 2, to accelerate convergence of SA, we employ inverse temperature scheduling with logarithmic function from $\beta_{\mathrm{init}}\coloneqq 0.9\,\beta_{\mathrm{nishimori}}$ to $\beta_{\mathrm{final}}\coloneqq\beta_{\mathrm{nishimori}}$
-$\beta_{i}\coloneqq\beta_{\mathrm{init}}\,\left(1+r\,\log{i}\right)\quad\text{for $i=1,\dots,N_{\mathrm{\beta}}$}$ (21)
+In Line 10 of Algorithm 3, we update the configuration $C$ by applying a randomly chosen stabilizer generator $G_{f}$.
+[^4] We formulate error configurations by bits, instead of Ising spins.
+Hence, multiplying configurations $C\mapsto C\,G_{f}$ is implemented as bitwise XOR (addition modulo two) on the two-bit representation per qubit.
+This update keeps the syndrome $S$ invariant.
+For simulated annealing (SA), we introduce a tempered family of unnormalized weights
+$\pi_{\beta}(C)\coloneqq e^{-\beta\,\mathcal{H}(C)},$ (22)
+where $\beta>0$ is treated as a tunable inverse temperature.
+In particular, the unnormalized Boltzmann weight corresponding to the physical error rate is recovered at $\beta=\beta_{\mathrm{N}}$, i.e., $\pi(C)=\pi_{\beta_{\mathrm{N}}}(C)$, which follows from Eq.
+(11).
+Therefore, $\beta_{\mathrm{N}}$ provides a natural fixed-temperature choice for MCMC-based decoding [nishimori_temperature, nishimori_book, Chubb_stat_model].
+In Algorithm 2, to accelerate convergence of SA, we employ a logarithmic schedule for the inverse temperature, increasing $\beta$ from $\beta_{\mathrm{init}}\coloneqq 0.9\,\beta_{\mathrm{N}}$ to $\beta_{\mathrm{final}}\coloneqq\beta_{\mathrm{N}}$
+$\beta_{i}\coloneqq\beta_{\mathrm{init}}\,\left(1+\gamma\,\log i\right)\quad\text{for $i=1,\dots,N_{\beta}$}$ (23)
 with
-$r\coloneqq\left.\left(\dfrac{\beta_{\mathrm{final}}}{\beta_{\mathrm{init}}}-1\right)\middle/\left(\log{N_{\mathrm{\beta}}}\right)\right.$. (22)
-For each inverse temperature $\beta$, we execute the Metropolis step $N_{\mathrm{stabilizer}}$ times, which is counted as one Monte Carlo step, according to the convention.
+$\gamma\coloneqq\begin{cases}\displaystyle\left.\left(\dfrac{\beta_{\mathrm{final}}}{\beta_{\mathrm{init}}}-1\right)\middle/\log N_{\beta}\right.&\text{for }N_{\beta}\geq 2,\\ 0&\text{for }N_{\beta}\in\{0,1\}.\end{cases}$ (24)
+For each inverse temperature $\beta$, we execute the Metropolis step $N_{\mathrm{stabilizer}}$ times, which is counted as one Monte Carlo sweep, according to the convention.
 Algorithm 1 Decoder using simulated annealing
-1: $S$ : Syndrome for error configuration
-2: Recovery configuration $C$
+1: $S$ : measured syndrome
+2: $C$ : recovery configuration, i.e., $\partial{C}=S$
 3: function SA_decoder($S$)
-4:   $T\leftarrow$ greedy_matching($S$) $\triangleright$ pure error
-5:   for all $P\in\{I,X,Z,Y\}$ do
-6:    $C_{L_{P}}\coloneqq T\,L_{P}$ $\triangleright$ Initial configuration for SA
+4:   $R\leftarrow$ greedy_matching($S$)
+5:   for all $P\in\{I,X,Y,Z\}$ do
+6:    $C_{L_{P}}\coloneqq R\,L_{P}$ $\triangleright$ Initial configuration for SA
 7:    for $i=1$ to $N_{\mathrm{SA}}$ do
-8:      $E^{(i)}_{L_{P}}\leftarrow$ simulated_annealing($C_{L_{P}}$)
-9:    $E_{L_{P}}\coloneqq\min\limits_{i=1,\dots,N_{\mathrm{SA}}}{E^{(i)}_{L_{P}}}$
+8:      $E^{(i)}_{L_{P}}\leftarrow$ simulated_annealing(copy($C_{L_{P}}$))
+9:    $E_{L_{P}}\coloneqq\min\limits_{i=1,\dots,N_{\mathrm{SA}}}{E^{(i)}_{L_{P}}} \triangleright$ estimate of $E^{\min}_{L_{P}}$
 10:
 11:
-12:   $P^{*}\coloneqq\operatorname*{\arg\min}\limits_{P\in\{I,X,Y,Z\}}{E_{L_{P}}}$ $\triangleright$ Eq. (20) 13:
-14:   return $C_{L_{P^{*}}}$
+12:   $P^{\ast}\coloneqq\operatorname*{\arg\min}\limits_{P\in\{I,X,Y,Z\}}{E_{L_{P}}}$ $\triangleright$ Eq. (21) 13:
+14:   return $C_{L_{P^{\ast}}}$ $\triangleright$ defined as $R\,L_{P^{\ast}}$ in Line 6
 Algorithm 2 Simulated annealing
 1: $C$ : error configuration
-2: Ground state energy (or lower excited state energy close to the ground state energy)
+2: $E_{\mathrm{min}}$ : an estimate of the minimum energy encountered during annealing
 3: function simulated_annealing($C$)
 4:   $E\leftarrow\mathcal{H}(C)$ $\triangleright$ Eq. (10)
-5:   $E_{\mathrm{min}}\leftarrow E$
-6:   $r\coloneqq\left.\left(\dfrac{\beta_{\mathrm{final}}}{\beta_{\mathrm{init}}}-1\right)\middle/\left(\log{N_{\mathrm{\beta}}}\right)\right$. $\triangleright$ Eq. (22) 7:
-8:   for $i=1$ to $N_{\mathrm{\beta}}$ do
-9:    $\beta\leftarrow\beta_{\mathrm{init}}\,\left(1+r\,\log{i}\right)$ $\triangleright$ Eq. (21)
-10:    for $N_{\mathrm{stabilizer}}$ times do
-11:      $\Delta E\leftarrow$ metropolis_Step($\beta$,  $C$)
-12:      $E\leftarrow E+\Delta E$
-13:      $E_{\mathrm{min}}\leftarrow\min\{E_{\mathrm{min}},\,E\}$ $\triangleright$ Update min energy
-14:
-15:   return $E_{\mathrm{min}}$
-Algorithm 3 Metropolis method which acts a stabilizer generator
-1: $\beta$ : inverse temperature,  $C$ : Error configuration
-2: Difference of energy of error configurations
-3: function Metropolis_Step($\beta$,  $C$)
-4:   Select a stabilizer generator $G_{f}$ randomly.
-5:   $\Delta E\leftarrow\mathcal{H}(C\,G_{f})-\mathcal{H}(C)$ 6:
-7:   Generate a uniform random probability $r\in[0,1]$.
-8:   if $\Delta E<0$ or $r<e^{-\beta\,\Delta E}$ then
-9:    $C\leftarrow C\,G_{f}$ $\triangleright$ Flips spins
-10:    return $\Delta E$
-11:   else
-12:    return 0
+5:
+$E_{\mathrm{min}}\leftarrow E$
+6:
+7:   for $i=1$ to $N_{\beta}$ do
+8:    $\beta\leftarrow\beta_{\mathrm{init}}\,\left(1+\gamma\,\log{i}\right)$ $\triangleright$ Eq. (23)
+9:    repeat $N_{\mathrm{stabilizer}}$ times do $\triangleright$ one Monte Carlo sweep
+10:      $\Delta E\leftarrow$ Metropolis_step($\beta$,  $C$)
+11:      $E\leftarrow E+\Delta E$
+12:      $E_{\mathrm{min}}\leftarrow\min\{E_{\mathrm{min}},\,E\}$ $\triangleright$ Update min energy
+13:
+14:   return $E_{\mathrm{min}}$
+Algorithm 3 Metropolis method that applies a stabilizer generator
+1:Input: $\beta$ : inverse temperature
+2:Input: $C$ : error configuration (modified in place)
+3: Energy difference of error configurations
+4: function Metropolis_step($\beta$,  $C$)
+5:   Select a stabilizer generator $G_{f}$ uniformly at random.
+6:   $\Delta E\leftarrow\mathcal{H}(C\,G_{f})-\mathcal{H}(C)$ 7:
+8:   Draw $r\sim\mathrm{Uniform}(0,1)$.
+9:   if $r<\min\{1,\exp(-\beta\,\Delta E)\}$ then
+10:    $C\leftarrow C\,G_{f}$ $\triangleright$ Applies $G_{f}$ (by bit flips)
+11:    return $\Delta E$
+12:   else
+13:    return 0
 IV.2 The MWPM decoder
 IV.2.1 The MWPM decoding algorithm
-We briefly review the algorithm of the MWPM decoder for the XZZX code [9].
-The MWPM decoder is formulated as an algorithm on the decoding graph $(\mathcal{V}_{\bullet},\mathcal{E}_{\bullet})$, introduced in Sec.
-III.
-For $u,v\in S_{\bullet}$, we define the weighted distance $w(u,v)$ between the vertices $u=(u_{x},u_{y})$ and $v=(v_{x},v_{y})$ as the shorter of the following two path lengths:
-  • The weighted Manhattan distance
-    $\alpha_{z}\lvert u_{x}-v_{x}\rvert\,+\,\alpha_{x}\lvert u_{y}-v_{y}\rvert$, (23)
-  • The length of ‘path via a virtual boundary vertex’ [35].
-For $u,v\in S_{\mathrm{gray}}$, this path consists of only horizontal edges, which connect $u$ and the left or right boundary and connect $v$ and the left or right boundary.
-The length of the path equals to
-[…]
-    For $u,v\in S_{\mathrm{white}}$, this path consists of only vertical edges, which connect $u$ and the top or bottom boundary and connect $v$ and the top or bottom boundary.
-The length of the path equals to
-[…]
+We briefly review the algorithm of the MWPM decoder for the XZZX code [xzzx_nature].
+The MWPM decoder is formulated on the decoding graph $(\mathcal{V}_{\bullet},\mathcal{E}_{\bullet})$, introduced in Sec. III. We index the vertices by integer coordinates as
+$\mathcal{V}_{\mathrm{white}}$ $=\{(i,j)\mid i\in\{1,\dots,d\},\;j\in\{1,\dots,d-1\}\},$
+$\mathcal{V}_{\mathrm{gray}}$ $=\{(i,j)\mid i\in\{1,\dots,d-1\},\;j\in\{1,\dots,d\}\}$
+where $i$ and $j$ denote the horizontal and vertical coordinates, respectively, and $(1,1)$ is the lower-left corner.
+For vertices $u,v\in S_{\bullet}$, we define the weight $w(u,v)$ between $u=(u_{x},u_{y})$ and $v=(v_{x},v_{y})$
+$w(u,v)\coloneqq\min\left\{w_{\mathrm{int}}(u,v),\,w_{\mathrm{bnd}}(u,v)\right\}.$ (25)
+Here, $w_{\mathrm{int}}(u,v)$ is the weighted Manhattan distance [xzzx_nature]
+$w_{\mathrm{int}}(u,v)\coloneqq\tilde{\alpha}_{z}\lvert u_{x}-v_{x}\rvert\,+\,\tilde{\alpha}_{x}\lvert u_{y}-v_{y}\rvert$, (26)
+with
+$\tilde{\alpha}_{\mu}\coloneqq-\log\!{\left(\frac{\tilde{p}_{\mu}}{1-p}\right)}\quad\text{for $\mu\in\{x,z\}$}$ (27)
+where we defined the probabilities $\tilde{p}_{x}\coloneqq p_{x}+p_{y}$ and $\tilde{p}_{z}\coloneqq p_{z}+p_{y}$.
+Note that MWPM is invariant under multiplying all edge weights by a positive constant.
+Therefore, we may freely rescale the weights without changing the matching solution.
+The alternative weight $w_{\mathrm{bnd}}(u,v)$ is the weight of connecting both vertices to the nearest boundary [MWPM_complexity]:
+  • For $u,v\in S_{\mathrm{white}}$, we define
+    $\begin{split}w_{\mathrm{bnd}}(u,v)\coloneqq\tilde{\alpha}_{x}(&\min\{\operatorname{dist}_{\mathrm{B}}(u),\operatorname{dist}_{\mathrm{T}}(u)\}\,+\\ &\min\{\operatorname{dist}_{\mathrm{B}}(v),\operatorname{dist}_{\mathrm{T}}(v)\}\,),\end{split}$ (28)
+    where $\operatorname{dist}_{\mathrm{B}}(u)\coloneqq u_{y}$ and $\operatorname{dist}_{\mathrm{T}}(u)\coloneqq d-u_{y}$ are the vertical distances from the vertex $u$ to the bottom and top boundaries, respectively, and similarly for $v$.
+  • For $u,v\in S_{\mathrm{gray}}$, we define
+    $\begin{split}w_{\mathrm{bnd}}(u,v)\coloneqq\tilde{\alpha}_{z}(&\min\{\operatorname{dist}_{\mathrm{L}}(u),\operatorname{dist}_{\mathrm{R}}(u)\}\,+\\ &\min\{\operatorname{dist}_{\mathrm{L}}(v),\operatorname{dist}_{\mathrm{R}}(v)\}\,),\end{split}$ (29)
+    where $\operatorname{dist}_{\mathrm{L}}(u)\coloneqq u_{x}$ and $\operatorname{dist}_{\mathrm{R}}(u)\coloneqq d-u_{x}$ are the horizontal distances from the vertex $u$ to the left and right boundaries, respectively, and similarly for $v$.
 We present the algorithm of the MWPM decoder in Algorithm 4.
-In Line 4, if the number of vertices $|S|$ is odd, to enable a perfect matching, the virtual boundary vertex is appended to the syndrome $S$ so that $|S|$ is an even number [35].
-In Line 6, we create the list of all possible candidates for pairing
-$K\coloneqq\{\text{unordered pair }(u,v)\,|\,u,v\in S_{\bullet},\,u\neq v\}.$ (26)
-In Line 8, the MWPM_graph function finds a minimum-weight perfect matching (MWPM) [36], i.e., a perfect matching $L\subset K$ such that the total distance $\sum_{(u,v)\in L}w(u,v)$ is minimum.
-For our implementation, we employ the C++ library, Blossom V [37].
+In Line 4, if the number of vertices $|S_{\bullet}|$ is odd, a virtual boundary vertex $b$ is appended to the syndrome $S_{\bullet}$ so that a perfect matching exists (i.e., $|S_{\bullet}|$ becomes even) [MWPM_complexity].
+In this case, we extend the weight function $w(\cdot,\cdot)$ by defining, for $u\in S_{\bullet}$,
+$w(u,b)\coloneqq\begin{cases}\tilde{\alpha}_{x}\min\{\operatorname{dist}_{\mathrm{B}}(u),\operatorname{dist}_{\mathrm{T}}(u)\},&\text{for }\bullet=\mathrm{white},\\ \tilde{\alpha}_{z}\min\{\operatorname{dist}_{\mathrm{L}}(u),\operatorname{dist}_{\mathrm{R}}(u)\},&\text{for }\bullet=\mathrm{gray}.\end{cases}$ (30)
+Note that $w_{\mathrm{bnd}}(u,v)=w(u,b)+w(v,b)$, corresponding to pairing both the defects $u,v$ with the boundary $b$.
+In Line 6, we create the set of (unordered) pairs
+$K\coloneqq\{\{u,v\}\,\mid\,u,v\in S_{\bullet},\,u\neq v\},$ (31)
+which are all possible candidates for pairing.
+In Line 8, the MWPM_graph function finds a minimum-weight perfect matching (MWPM) [Edmonds], i.e., a perfect matching $L\subset K$ such that the total distance $\sum_{\{u,v\}\in L}w(u,v)$ is minimum.
+For our implementation, we employ the C++ library, Blossom V [BlossomV].
 Algorithm 4 The MWPM decoder
-1: $S_{\bullet}$ : Syndrome with $\bullet\in\{\mathrm{white},\mathrm{gray}\}$
-2: Recovery chain $C_{\bullet}$ consistent with $S_{\bullet}$
+1: $S_{\bullet}$ : Measured syndrome with $\bullet\in\{\mathrm{white},\mathrm{gray}\}$
+2: $C_{\bullet}$ : Recovery chain, i.e., $\partial{C_{\bullet}}=S_{\bullet}$
 3: function MWPM_chain($S_{\bullet}$)
 4:   if $|S_{\bullet}|$ is odd then
-5:    Append ‘the virtual boundary vertex’ to $S_{\bullet}$
-6:   $K\coloneqq\{\text{unordered pair }(u,v)\,|\,u,v\in S_{\bullet},\,u\neq v\}$ $\triangleright$ Eq. (26) 7:
+5:    Append the virtual boundary vertex $b$ to $S_{\bullet}$
+6:   $K\coloneqq\{\{u,v\}\,\mid\,u,v\in S_{\bullet},\,u\neq v\}$ $\triangleright$ Eq. (31) 7:
 8:   $L\leftarrow$ MWPM_graph($K$) $\triangleright$ Utilize Blossom V 9:
 10:   $C_{\bullet}\leftarrow\emptyset$
-11:   for all $(u,v)\in L$ do
-12:    Add edges which connect $u$ and $v$ to $C_{\bullet}$.
+11:   for all $\{u,v\}\in L$ do
+12:    Add to $C_{\bullet}$ the edges along a path connecting $u$ and $v$ with the weight $w(u,v)$.
 13:
 14:   return $C_{\bullet}$
-1: $S\coloneqq(S_{\mathrm{white}},\,S_{\mathrm{gray}})$ : Syndrome
-2: Recovery configuration $C$ consistent with $S$
+1: $S\coloneqq(S_{\mathrm{white}},\,S_{\mathrm{gray}})$ : Measured syndrome
+2: $C$ : Recovery configuration, i.e., $\partial{C}=S$
 3: function MWPM($S$)
@@
 8:   return $C$
-In Line 12, for each matching pair $(u,v)\in L$, we connect the vertices $u$ and $v$ with a path which consists of edges, i.e., applying $Z$ ($X$) operators to the qubits located at the centers of horizontal (vertical) edges, respectively.
+In Line 12, for each matched pair $\{u,v\}\in L$, we connect the vertices $u$ and $v$ with a path which consists of edges, i.e., applying $Z$ and $X$ operators to the qubits located at the centers of horizontal and vertical edges, respectively.
+This path has weight $w(u,v)$, i.e., it is chosen to attain the distance defined by Eqs.
+(26), (28), (29), or (30).
 The set of these edges forms a recovery chain $C_{\bullet}$.
-The MWPM function invokes the MWPM_CHAIN function twice, i.e., for the syndromes $S_{\mathrm{white}}$ and $S_{\mathrm{gray}}$ on the white gray and faces, respectively.
-Note that, the syndromes on only faces of the same color may be connected.
-The MWPM algorithm is too complicated to implement in a hardware system such as FPGA. Thus, we propose a simple variant algorithm in Sec. IV.3.
-IV.2.2 Interpreting the MWPM decoder as an energy minimization problem
-Let us interpret the MWPM decoder as energy minimization problem, then compare it with our SA decoder.
-Since the perfect matching is equivalent to the syndrome consistency $\partial{C_{\bullet}}=S$, the total distance of such an error chain $C_{\bullet}$ is defined as
-$w(C_{\bullet})\coloneqq\sum_{(u,v)\in L_{0}}w(u,v)$ (27)
-where $L_{0}$ is a perfect matching. Then, the MWPM decoding algorithm is equivalent to find a recovery error chain $C_{\bullet}$:
-$\arg\min_{C_{\bullet}}\left\{w(C_{\bullet})\,\middle|\,\partial{C_{\bullet}}=S\right\}.$ (28)
-Since the weighted distance $w(u,v)$ is defined by Eqs. (23),(24) or (25), Eq. (27) is expressed as
-$w(C_{\bullet})=\alpha_{z}\,n_{\mathrm{hor}}\,+\,\alpha_{x}\,n_{\mathrm{ver}}$ (29)
-where $n_{\mathrm{hor}}$ ($n_{\mathrm{ver}}$) is the number of errors which reside on horizontal (vertical) edges in the error chain $C_{\bullet}$, respectively.
-In Eq.
-(29), the weight coefficients $\alpha_{x},\alpha_{z}$ can be calculated by Eq.
-(9) from the error probabilities $p_{x},p_{z}$ assuming we know.
-Let us compare the total distance in Eq.
+The MWPM function invokes the MWPM_chain function twice, i.e., for the syndromes $S_{\mathrm{white}}$ and $S_{\mathrm{gray}}$ on the white and gray faces, respectively.
+Note that only syndromes on faces of the same color may be connected.
+Implementing the MWPM decoder efficiently on hardware such as FPGAs is challenging due to its algorithmic complexity and memory/communication requirements.
+Thus, we propose a simple variant algorithm in Sec.
+IV.3.
+IV.2.2 The MWPM decoder vs. our SA decoder as a minimization problem
+Given a perfect matching $L\subset K$, we define its total weight as
+$w(L)\coloneqq\sum_{\{u,v\}\in L}w(u,v).$ (32)
+The MWPM decoder finds a minimum-weight perfect matching $L^{\ast}$:
+$L^{\ast}\in\arg\min_{L}\left\{w(L)\,\middle|\,L\text{ is a perfect matching on }S_{\bullet}\right\}.$
+From a perfect matching $L$, we can construct a recovery chain $C_{\bullet}(L)$, i.e., $\partial{C_{\bullet}(L)}=S_{\bullet}$, by connecting each matched pair $\{u,v\}\in L$ with a path of weight w(u,v)w(u,v) [^5] Note that $C_{\bullet}(L)$ is not unique in general, due to degeneracies in the minimum-weight paths..
+Since the weighted distance $w(u,v)$ is defined by Eqs.
+(26),(28), (29), or (30), Eq.
+(32) is expressed as
+$w(L)=\tilde{\alpha}_{z}\,n_{\mathrm{hor}}\,+\,\tilde{\alpha}_{x}\,n_{\mathrm{ver}}$ (33)
+where $n_{\mathrm{hor}}$ and $n_{\mathrm{ver}}$ are the numbers of occupied horizontal and vertical edges in the error chain $C_{\bullet}(L)$, i.e., the numbers of $Z$- and $X$-type components, respectively.
+A $Y$ error contributes to both $n_{\mathrm{hor}}$ and $n_{\mathrm{ver}}$.
+We now compare the total distance in Eq.
 (33) and the energy defined in Eq.
 (10).
-Since Eq.
-(29) does not contain the probability $p_{y}$, the MWPM decoder cannot properly consider $Y$ errors unless $p_{x}:p_{y}:p_{z}=1:2:1$, and therefore is not optimal.
-On the other hand, Eq.
-(10) contains the probability $p_{y}$, and hence our SA decoder is able to attain optimal accuracy.
-In Sec.
-V, we execute the MWPM decoder to evaluate the accuracy of our SA decoders for comparison.
+In the MWPM approach for the XZZX code, one solves two separate matching problems on the white and gray decoding graphs.
+This effectively decouples the inference of the $X$- and $Z$-type components and does not model the fact that a $Y$ error simultaneously contributes to both components on the same data qubit; hence it cannot capture $Y$-induced correlations.
+Consequently, when $p_{y}$ is non-negligible, this MWPM approach is expected to be suboptimal compared to our SA decoder, which explicitly incorporates $p_{y}$ via the $Y$-error weight (through $\alpha_{y}$) in the energy $\mathcal{H}(C)$ in Eq.
+(10).
 IV.3 The greedy matching decoder
@@
 IV.2.1, the MWPM decoder utilizes the MWPM graph algorithm, i.e., the perfect matching giving the minimum of the total distance.
-Exchanging this ‘global’ minimality for a low computational cost and simple algorithm yields a greedy version of the perfect matching algorithm.
-In this subsection, we elaborate on our greedy matching algorithm.
-We employ this algorithm to construct a pure error $T(S)$, set as an initial configuration for our SA decoder in Algorithm 1.
-We give our algorithm in Algorithm 5.
-In Line 10, the GREEDY_MATCHING_CHAIN function matches the closest pair $(u,v)\in K$ with respect to $w(u,v)$, and connects the vertices $u$ and $v$.
-In Line 12, it excludes all the pairs whose vertex is $u$ or $v$ from the candidate list $K$.
+By replacing the global minimization of MWPM with a local greedy matching rule, we obtain a simple and low computational cost decoder.
+We employ the greedy matching decoding algorithm to construct a reference configuration $R(S)$, set as an initial configuration for our SA decoder in Algorithm 1.
+We give our greedy matching decoding algorithm in Algorithm 5.
+In Line 10, the greedy_matching_chain function matches the closest pair $\{u,v\}\in K$ with respect to $w(u,v)$, and connects the vertices $u$ and $v$.
+Line 12 removes all pairs that contain $u$ or $v$ from the candidate list $K$ of vertex pairs.
 This algorithm thus enables us to monotonically reduce the size of the list $K$.
-We efficiently implement $K$ using priority queue [38, §6.5].
-In the priority queue with $n$ elements (e.g., a binary heap), each of insertion and extract-min operations takes $\mathcal{O}(\log n)$; therefore, performing the operations for $n$ elements has time complexity $\mathcal{O}(n\log n)$.
-For our greedy matching decoder, the size of the priority queue is $n=\mathcal{O}(|S_{\bullet}|^{2})$, where $|S_{\bullet}|$ denotes the number of incorrect syndromes.
-Thus, the time complexity of our greedy matching decoder is $\mathcal{O}(|S_{\bullet}|^{2}\log{|S_{\bullet}|})$.
+We efficiently implement $K$ using a priority queue [algorithm_Cormen, §6.5].
+In the priority queue with $m$ elements (e.g., a binary heap), each of insertion and extract-min operations takes $\mathcal{O}(\log m)$; therefore, performing the operations for $m$ elements has time complexity $\mathcal{O}(m\log m)$.
+For our greedy matching decoder, the size of the priority queue is $m=\mathcal{O}(|S_{\bullet}|^{2})$, where $|S_{\bullet}|$ denotes the number of syndrome defects.
+Thus, the worst-case time complexity is $\mathcal{O}(|S_{\bullet}|^{2}\log|S_{\bullet}|)$ for each color.
+Since $|S_{\bullet}|\leq|\mathcal{V}_{\bullet}|=\mathcal{O}(d^{2})$, this yields the worst-case bound $\mathcal{O}(d^{4}\log d)$.
 Algorithm 5 The greedy matching decoder
-1: $S_{\bullet}$ : Syndrome with $\bullet\in\{\mathrm{white},\mathrm{gray}\}$
-2: Recovery chain $C_{\bullet}$ consistent with $S_{\bullet}$
+1: $S_{\bullet}$ : Measured syndrome with $\bullet\in\{\mathrm{white},\mathrm{gray}\}$
+2: $C_{\bullet}$ : Recovery chain, i.e., $\partial{C_{\bullet}}=S_{\bullet}$
 3: function greedy_matching_chain($S_{\bullet}$)
 4:   if $|S_{\bullet}|$ is odd then
-5:    Append ‘the virtual boundary vertex’ to $S_{\bullet}$
-6:   $K\coloneqq\{\text{unordered pair }(u,v)\,|\,u,v\in S_{\bullet},\,u\neq v\}$ $\triangleright$ Eq. (26)
+5:    Append the virtual boundary vertex $b$ to $S_{\bullet}$
+6:   $K\coloneqq\{\{u,v\}\,|\,u,v\in S_{\bullet},\,u\neq v\}$ $\triangleright$ Eq. (31)
 7:   $C_{\bullet}\leftarrow\emptyset$ 8:
 9:   while $K$ is not empty do
-10:    Pick a pair $(u,v)\in K$ such that $w(u,v)$ is minimum.
-11:    Add edges which connect $u$ and $v$ to $C_{\bullet}$.
-12:    Exclude all the pairs whose vertex is $u$ or $v$ from the list $K$.
+10:    Pick a pair $\{u,v\}\in K$ such that $w(u,v)$ is minimum.
+11:    Add to $C_{\bullet}$ the edges along a path connecting $u$ and $v$ with the weight $w(u,v)$.
+12:    Remove from $K$ all pairs $\{x,y\}$ such that $x\in\{u,v\}$ or $y\in\{u,v\}$.
 13:
 14:   return $C_{\bullet}$
-1: $S\coloneqq(S_{\mathrm{white}},\,S_{\mathrm{gray}})$ : Syndrome
-2: Recovery configuration $C$ consistent with $S$
+1: $S\coloneqq(S_{\mathrm{white}},\,S_{\mathrm{gray}})$ : Measured syndrome
+2: $C$ : Recovery configuration, i.e., $\partial{C}=S$
 3: function greedy_matching($S$)
@@
 8:   return $C$
-As well as the MWPM decoder, since our greedy matching decoder employs the weighted distance $w(u,v)$ which is calculated from the probabilities $p_{z}$ and $p_{x}$, it cannot take into account the probability $p_{y}$.
-IV.4 Our SA decoder which starts with different pure errors
+Like the MWPM decoder, our greedy matching decoder employs the weighted distance $w(u,v)$, which is calculated from $p_{z}$ and $p_{x}$, and therefore cannot take $p_{y}$ into account.
+IV.4 Our SA decoder which starts with different reference configurations
 In Sec. IV.1, we described our SA decoder in Algorithm 1, which performs SA multiple times to achieve energy as low as possible.
-To improve efficiency, rather than using an identical pure error repeatedly, it is preferable to use a different pure error for each SA run.
-To this end, we modify our greedy matching algorithm.
-Specifically, when multiple pairs of vertices have the same distance, their matching order is randomized using pseudo-random numbers.
+To improve efficiency, rather than using an identical reference configuration repeatedly, it is preferable to use a different reference configuration for each SA run.
+To this end, we modify our greedy matching algorithm by randomizing the tie-breaking among equal-weight pairs.
+Specifically, we randomize tie-breaking among equal-weight pairs using pseudo-random numbers.
 This modified algorithm is presented in Algorithm 6.
-In Algorithm 6, we create the candidate list $K_{w}$, defined as the set of pairs which give the distance $w$:
-$K_{w}\coloneqq\{(u,v)\in K\,|\,w(u,v)=w\}.$ (30)
-This newly introduced randomness may produce a different pure error for every invocation, while still remaining consistent with the given syndrome $S$.
-However, such pure errors differ not only in their stabilizer components but also in their logical components in the decomposition of Eq.
+In Algorithm 6, we form the subset
+$K_{\min}\coloneqq\{\{u,v\}\in K\,\mid\,w(u,v)=w_{\min}\},$ (34)
+i.e., the set of candidate pairs that attain the current minimum weight $w_{\min}$.
+This newly introduced randomness may produce a different reference configuration for every invocation, while still satisfying $\partial{C}=S$.
+However, two such reference configurations can differ by both a stabilizer component $G$ and a logical operator component $L_{P}$ in the decomposition of Eq.
 (14).
-Therefore, we need to transform a logical operator to the one based on a ‘standard’ pure error, as follows.
-We present SA decoder which utilizes Algorithm 6 in Algorithm 7. In Line 5, for the first SA, we construct a pure error $T_{1}(S)$, which is used as ‘standard’.
-Let us denote by $T(S)$ a pure error constructed from the second SA onward. By applying Eq. (13), $T_{1}(S)$ can be decomposed as
-$T_{1}(S)=T(S)\,L_{Q}\,G^{\prime}$ (31)
-for $Q\in\{I,X,Z,Y\}$ and $G^{\prime}\in\mathcal{G}$.
-The logical operator $L_{Q}$ expresses ‘difference’ of equivalence classes of logical operators based on $T(S)$ and $T_{1}(S)$, and is obtained by $L_{Q}]=[T(S)\,T_{1}(S)]$.
-From Eq. (31), it follows that
-$\operatorname{Rep}_{T(S)}(L_{P})=\operatorname{Rep}_{T_{1}(S)}(L_{QP}),$
-which implies that “the logical operator $L_{P}$ based on T(S)T(S)” equals to “the logical operator $L_{QP}$ based on $T_{1}(S)$”.
-Thus, for each $P\in\{I,X,Z,Y\}$, we store the calculated energy in $E^{(i)}_{L_{QP}}$ in Line 16.
-It should be noted that, in the for loops in Lines 11 and 14 in Algorithm 7, each iteration is an independent task, and hence these tasks can obviously be parallelized.
-In addition, Line 13 can be parallelized with Lines 1516.
-Algorithm 6 The greedy matching decoder which may produce a different pure error
-1: $S_{\bullet}$ : Syndrome with $\bullet\in\{\mathrm{white},\mathrm{gray}\}$
-2: Recovery chain $C_{\bullet}$ consistent with $S_{\bullet}$
+Therefore, to compare energies across runs consistently, we relabel the logical operator cosets obtained in each run as those defined with respect to a fixed “standard” reference configuration.
+We present a second SA decoder that utilizes Algorithm 6 in Algorithm 7. In Line 4, we construct a reference configuration $R_{1}(S)\in\Omega_{S}$, which we use as the fixed standard reference.
+Let $R_{i}(S)\in\Omega_{S}$ ($i\geq 2$) denote the reference configuration produced in the $i$th run. Using Eq. (14), $R_{i}(S)$ can be decomposed as
+$R_{i}(S)=R_{1}(S)\,L_{Q}\,G^{\prime}$ (35)
+for some $Q\in\{I,X,Y,Z\}$ and some $G^{\prime}\in\mathcal{G}$.
+Since $\partial(R_{i}(S)\,R_{1}(S))=\emptyset$, the equivalence class $R_{i}(S)\,R_{1}(S)]\in\overline{\mathcal{P}}/\mathcal{G}$ is well-defined and we obtain $R_{i}(S)\,R_{1}(S)]=[L_{Q}]$.
+The Pauli operator $Q$ can be determined by checking its commutation relations with the logical operators $L_{X}$ and $L_{Z}$.
+Since stabilizers commute with logical operators, Eq. (35) implies the correspondence
+$\operatorname{\mathcal{C}}_{R_{i}(S)}(L_{P})=\operatorname{\mathcal{C}}_{R_{1}(S)}(L_{QP}).$ (36)
+where $QP$ denotes the Pauli multiplication in $\{I,X,Y,Z\}$ (up to global phase), consistent with our convention of ignoring global phases.
+Eq.
+(36) means that “the coset of $L_{P}$ with respect to $R_{i}(S)$” equals “the coset of $L_{QP}$ with respect to $R_{1}(S)$”.
+Thus, for each $P\in\{I,X,Y,Z\}$, we store the calculated energy in $E^{(i)}_{L_{QP}}$ in Line 14.
+Note that, in the for loops in Lines 9 and 12 in Algorithm 7, each iteration is an independent task, and hence these tasks can trivially be parallelized.
+In addition, Line 11 can be parallelized with Lines 1314.
+Algorithm 6 The greedy matching decoder which may produce a different reference configuration
+1: $S_{\bullet}$ : Measured syndrome with $\bullet\in\{\mathrm{white},\mathrm{gray}\}$
+2: $C_{\bullet}$ : Recovery chain, i.e., $\partial{C_{\bullet}}=S_{\bullet}$
 3: function greedy_matching_diff_chain($S_{\bullet}$)
 4:   if $|S_{\bullet}|$ is odd then
-5:    Append ‘the virtual boundary vertex’ to $S_{\bullet}$
-6:   $K\coloneqq\{\text{unordered pair }(u,v)\,|\,u,v\in S_{\bullet},\,u\neq v\}$ $\triangleright$ Eq. (26)
+5:    Append the virtual boundary vertex $b$ to $S_{\bullet}$
+6:   $K\coloneqq\{\{u,v\}\,|\,u,v\in S_{\bullet},\,u\neq v\}$ $\triangleright$ Eq. (31)
 7:
@@
 9:   while $K$ is not empty do
-10:    $w\coloneqq\min\{w(u,v)\,|\,(u,v)\in K\}$
-11:    $K_{w}\coloneqq\{(u,v)\in K\,|\,w(u,v)=w\}$ $\triangleright$ Eq. (30)
-12:    while $K_{w}$ is not empty do
-13:      Pick a pair $(u,v)\in K_{w}$ randomly.
-14:      Add edges which connect $u$ and $v$ to $C_{\bullet}$.
-15:      Exclude all the pairs whose vertex is $u$ or $v$ from the lists $K_{w}$ and $K$.
+10:    $w_{\min}\coloneqq\min\{w(u,v)\,|\,\{u,v\}\in K\}$
+11:    $K_{\min}\coloneqq\{\{u,v\}\in K\,|\,w(u,v)=w_{\min}\}$ $\triangleright$ Eq. (34)
+12:    while $K_{\min}$ is not empty do
+13:      Pick a pair $\{u,v\}\in K_{\min}$ uniformly at random.
+14:      Add to $C_{\bullet}$ the edges along a path connecting $u$ and $v$ with the weight $w(u,v)$.
+15:      Remove from $K$ and $K_{\min}$ all pairs $\{x,y\}$ such that $x\in\{u,v\}$ or $y\in\{u,v\}$.
 16:
 17:   return $C_{\bullet}$
-1: $S\coloneqq(S_{\mathrm{white}},\,S_{\mathrm{gray}})$ : Syndrome
-2: Recovery configuration $C$ consistent with $S$
+1: $S\coloneqq(S_{\mathrm{white}},\,S_{\mathrm{gray}})$ : Measured syndrome
+2: $C$ : Recovery configuration, i.e., $\partial{C}=S$
 3: function greedy_matching_different($S$)
@@
 8:   return $C$
-Algorithm 7 Decoder using simulated annealing which may start with different pure errors
-1: $S$ : Syndrome for error configuration
-2: Recovery configuration $C$
-3: function SA_Decoder_Different_pure_error($S$)
-4:   $\triangleright$  First SA   $\triangleleft$
-5:   $T_{1}\leftarrow$ greedy_matching_different($S$)
-6:   for all $P\in\{I,X,Z,Y\}$ do
-7:    $C^{(1)}_{L_{P}}\coloneqq T_{1}\,L_{P}$ $\triangleright$ Initial configuration for SA
+Algorithm 7 Decoder using simulated annealing which may start with different reference configurations
+1: $S$ : Measured syndrome
+2: $C$ : Recovery configuration, i.e., $\partial{C}=S$
+3: function SA_decoder_different_reference($S$)
+4:   $R_{1}\leftarrow$ greedy_matching_different($S$) $\triangleright$ standard reference configuration
+5:   for all $P\in\{I,X,Y,Z\}$ do
+6:    $C^{(1)}_{L_{P}}\coloneqq R_{1}\,L_{P}$ $\triangleright$ standard reference
 […]
-9:
-10:   $\triangleright$  Second and later SAs   $\triangleleft$
-11:   for $i=2$ to $N_{\mathrm{SA}}$ do
-12:    $T\leftarrow$ greedy_matching_different($S$)
-13:    Determine $Q\in\{I,X,Z,Y\}$ such that $L_{Q}]=[T\,T_{1}]$.
-14:    for all $P\in\{I,X,Z,Y\}$ do
-15:      $C_{L_{P}}\coloneqq T\,L_{\mathrm{P}}$ $\triangleright$ Initial configuration for SA
-16:      $E^{(i)}_{L_{QP}}\leftarrow$ simulated_annealing($C_{L_{P}}$)
-17:
-18:   for all $P\in\{I,X,Z,Y\}$ do
-19:    $E_{L_{P}}\coloneqq\min\limits_{i=1,\dots,N_{\mathrm{SA}}}{E^{(i)}_{L_{P}}}$
-20:
-21:   $P^{*}\coloneqq\operatorname*{\arg\min}\limits_{P\in\{I,X,Y,Z\}}{E_{L_{P}}}$ $\triangleright$ Eq. (20) 22: 23:
-24:   return $C^{(1)}_{L_{P^{*}}}$
-Like the MWPM decoder, our greedy matching decoder is MLE-type; the MLD variant is obtained by disabling SA (i.e., by setting $N_{\beta}=0$) in Algorithm 7.
-V Numerical result
+8:
+9:   for $i=2$ to $N_{\mathrm{SA}}$ do
+10:    $R_{i}\leftarrow$ greedy_matching_different($S$)
+11:    Determine $Q\in\{I,X,Y,Z\}$ such that $L_{Q}]=[R_{i}\,R_{1}]$. $\triangleright$ e.g., via commutation with $L_{X}$ and $L_{Z}$
+12:    for all $P\in\{I,X,Y,Z\}$ do
+13:      $C^{(i)}_{L_{P}}\coloneqq R_{i}\,L_{P}$ $\triangleright$ Initial configuration for SA
+[…]
+15:
+16:   for all $P\in\{I,X,Y,Z\}$ do
+17:    $E_{L_{P}}\coloneqq\min\limits_{i=1,\dots,N_{\mathrm{SA}}}{E^{(i)}_{L_{P}}} \triangleright$ estimate of $E^{\min}_{L_{P}}$
+18:
+19:   $P^{\ast}\coloneqq\operatorname*{\arg\min}\limits_{P\in\{I,X,Y,Z\}}{E_{L_{P}}}$ $\triangleright$ Eq. (21) 20: 21:
+22:   return $C^{(1)}_{L_{P^{\ast}}} \triangleright$ defined in Line 6
+V Numerical results
 We perform Monte Carlo simulations of our SA decoder under the setting described in Sec.
@@
 V.1 Simulation setting
-In the Monte Carlo simulation, we generate samples of actual error configurations $C$ by means of pseudo-random numbers, drawn from the physical error rate $p$ and the error bias $p_{x}:p_{y}:p_{z}$ we specify.
-Performing the stabilizer measurement on the actual error configurations $C$ yields the syndrome $S$.
-Then, we execute each decoder to obtain the recovery configuration $\bar{C}$.
-If the product $C\,\bar{C}$ belongs to the (non-trivial) logical operator class $L_{X}],[L_{Y}]$ or $L_{Z}]$, then the logical error occurs.
+In Monte Carlo simulations, we generate samples of true error configurations $\bar{C}$ by means of pseudo-random numbers, drawn from the physical error rate $p$ and the error bias $p_{x}:p_{y}:p_{z}$ we specify.
+Performing the stabilizer measurement on the true error configurations $\bar{C}$ yields the syndrome $S$.
+Then, we execute each decoder to obtain the recovery configuration $C$.
+If the product $\bar{C}\,C$ belongs to a nontrivial logical operator class, i.e., if $\bar{C}\,C]\in\{[L_{X}],[L_{Y}],[L_{Z}]\}$, then the logical error is said to occur.
 By repeating the aforementioned procedure $N_{\mathrm{sample}}$ times, we estimate the logical error rate $P_{\mathrm{L}}$.
 Except in Sec.
-V.4, we employ $N_{\mathrm{sample}}=10^{6}$ to demand sufficient statistical precision.
+V.4, we employ $N_{\mathrm{sample}}=10^{6}$ to ensure sufficient statistical precision.
 Except in Sec.
-V.3, we refer to Algorithm 7 as ‘our SA decoder’ and employ the number of inverse temperatures $N_{\beta}=100$.
-Although our SA decoder can be applied to any type of biased noise, we focus on $Y$-biased noise such that $X$ and $Z$ errors occur with the same probabilities $p_{x}=p_{z}$, which is the simplest case where the MWPM decoder cannot produce a correct recovery configuration.
-By the condition $p_{x}=p_{z}$, we can express as $p_{x}:p_{y}:p_{z}=1:\eta:1$ where $\eta$ indicates the $Y$-biased factor.
-Since the depolarizing noise is the case where $p_{x}=p_{y}=p_{z}=1/3$, it corresponds to $p_{x}:p_{y}:p_{z}=1:1:1$, i.e., $\eta=1$.
-The more $Y$-biased, the less accurate the MWPM decoder.
+V.3, we refer to Algorithm 7 as our SA decoder and employ the number of inverse temperatures $N_{\beta}=100$.
+Our SA decoder is applicable to arbitrary biased Pauli noise provided that the decoder is supplied with the error probabilities $p_{x},p_{y},p_{z}$ used to define $\mathcal{H}(C)$.
+We focus on $Y$-biased noise as a stress test for the MWPM decoder, which neglects $Y$-induced correlations.
+For simplicity, we consider the symmetric case $p_{x}=p_{z}$.
+Under the condition $p_{x}=p_{z}$, we can write the bias as $p_{x}:p_{y}:p_{z}=1:\eta:1$ where $\eta$ denotes the $Y$-biased factor.
+Since the depolarizing noise is the case where $p_{x}=p_{y}=p_{z}=p/3$, it corresponds to $p_{x}:p_{y}:p_{z}=1:1:1$, i.e., $\eta=1$.
 V.2 Logical error rates
-Let us compare the decoding accuracy of our SA decoder described in Algorithm 7 and the CPLEX decoder described in Sec.
-A.
+Let us compare the decoding accuracy of our SA decoder described in Algorithm 7 with the CPLEX decoder described in Appendix A.
+The CPLEX decoder serves as an optimal minimum-energy (MAP-configuration) decoder for the objective $\mathcal{H}(C)$ under the constraint $\partial C=S$.
 First, we examine how the logical error rate varies with the code distance $d$.
 Fig.
-3 depicts the logical error rates $P_{\mathrm{L}}$ under the depolarizing noise and for the code distances $d=5,7,9$.
+3 depicts the logical error rates $P_{\mathrm{L}}$ under the depolarizing noise for the code distances $d=5,7,9$.
 In Fig.
-3, for every $d$, our SA decoder achieves CPLEX’s rates, which is optimal among MLE-type decoders.
-In lower physical error rates $p$, our SA decoder outperforms the CPLEX decoder.
-This is because, unlike the CPLEX decoder, our SA decoder accounts for the degeneracy of error configurations by setting the final temperature to the Nishimori temperature $\beta_{\mathrm{nishimori}}$ in Eq.
-(12).
-For both the SA and CPLEX decoders, from the intersection point of these curves, we determine the threshold $p_{\mathrm{th}}\simeq 0.18$, which approximately equal to $p_{\mathrm{th}}=0.187$ obtained by the MPS decoder [9].
-Below the threshold $p_{\mathrm{th}}$, as $d$ increases, the logical error rate decreases, as expected.
-Figure 3: Logical error rates under the depolarizing noise and the code distance $d=5,7,9$.
+3, our SA decoder achieves the logical error rates of the CPLEX decoder within statistical uncertainty for all $d$ shown.
+For both the SA and CPLEX decoders, from the crossing of the finite-distance curves, we estimate a finite-size threshold $p_{\mathrm{th}}\simeq 0.18$.
+This estimate is close to the previously reported value $p_{\mathrm{th}}=0.187$ obtained using an MPS decoder [xzzx_nature].
+Below the threshold $p_{\mathrm{th}}$, the logical error rate decreases with increasing $d$, as expected.
+Figure 3: Logical error rates under the depolarizing noise for the code distances $d=5,7,9$.
 Our SA decoder achieves the logical error rates of the CPLEX decoder for all $d$.
-We employ $N_{\mathrm{SA}}=100$, which suffices to attain CPLEX’s accuracy.
+We employ $N_{\mathrm{SA}}=100$, which suffices to attain the logical error rates of the CPLEX decoder.
 The error bars indicate one standard deviation.
@@
 In Fig.
-4, for every $d$, our SA decoder achieves CPLEX’s rates, which are optimal.
-For both SA and CPLEX decoders, the intersection point of these curves lies around $p=0.2$, which we consider to be the threshold.
+4, for every $d$, our SA decoder achieves the logical error rates of the CPLEX decoder, which are optimal.
+For both SA and CPLEX decoders, the intersection point of these curves lies around $p\simeq 0.2$, which we consider to be a finite-size threshold estimate.
 Below this threshold, the logical error rate decreases as $d$ increases.
-Figure 4: Logical error rates under the $Y$-biased noise where $p_{x}:p_{y}:p_{z}=1:5:1$ and for the code distance $d=5,7,9$.
+Figure 4: Logical error rates under the $Y$-biased noise where $p_{x}:p_{y}:p_{z}=1:5:1$ for the code distances $d=5,7,9$.
 Our SA decoder achieves the logical error rates of the CPLEX decoder.
-We employ $N_{\mathrm{SA}}=100$, which suffices to attain CPLEX’s accuracy.
+We employ $N_{\mathrm{SA}}=100$, which suffices to attain the logical error rates of the CPLEX decoder.
 The error bars indicate one standard deviation.
@@
 In Fig.
-5, no intersection point is observed for $d=5,7,9$; hence, we infer that the threshold lies beyond the plotted range, i.e., $p_{\mathrm{th}}>0.2$.
+5, no intersection point for the CPLEX decoder is observed for $d=5,7,9$; hence, we expect that the threshold for the CPLEX decoder lies beyond the plotted range, i.e., $p_{\mathrm{th}}>0.2$.
 On the other hand, for $d=9$ in Fig.
-5, our SA decoder does not achieve CPLEX’s rate.
-The curves for $d=5,7,9$ do not intersect, due to the critical slowing down of the MCMC method.
-For our SA decoder, the curves for $d=7,9$ intersect around $p=0.11$, which we regarded as the threshold.
+5, our SA decoder does not achieve the logical error rates of the CPLEX decoder.
+The lack of a clear crossing for $d=5,7,9$ may be related to slow mixing of the MCMC method.
+For our SA decoder, the curves for $d=7,9$ intersect around $p=0.11$, which we take as an estimated finite-size threshold.
 When $p<0.11$, the logical error rate decreases as $d$ increases.
-Figure 5: Logical error rates under the $Y$-biased noise where $p_{x}:p_{y}:p_{z}=1:10:1$ and for the code distance $d=5,7,9$.
+Figure 5: Logical error rates under the $Y$-biased noise where $p_{x}:p_{y}:p_{z}=1:10:1$ for the code distances $d=5,7,9$.
 Our SA decoder cannot achieve the logical error rate of the CPLEX decoder for $d=9$, even by $N_{\mathrm{SA}}=100$.
 The error bars indicate one standard deviation.
-Next, we investigate the dependence of the logical error rates $P_{\mathrm{L}}$ on the number of SA, $N_{\mathrm{SA}}$ for fixed code distance $d=9$.
+Next, we investigate the dependence of the logical error rates $P_{\mathrm{L}}$ on the number of SA runs, $N_{\mathrm{SA}}$ for fixed code distance $d=9$.
 For comparison, we also calculate the rates by our greedy matching decoder, i.e., the greedy_matching function described in Algorithm 5 and the MWPM decoder i.e., the MWPM function described in Algorithm 4.
@@
 7 and 8 show the logical error rates under $Y$-biased noises where $p_{x}:p_{y}:p_{z}=1:5:1$ and $1:10:1$ for several values of $N_{\mathrm{SA}}$, respectively.
-Figure 6: Logical error rates under the depolarizing noise and $N_{\mathrm{SA}}=1,\dots,10$. We employ the code distance $d=9$. The error bars indicate one standard deviation.
-Figure 7: Logical error rates under $Y$-biased noise where $p_{x}:p_{y}:p_{z}=1:5:1$ and $N_{\mathrm{SA}}=1,\dots,100$.
-We employ the code distance $d=9$.
-The error bars indicate one standard deviation.
-Figure 8: Logical error rates under $Y$-biased noise where $p_{x}:p_{y}:p_{z}=1:10:1$ and $N_{\mathrm{SA}}=1,\dots,100$.
-We employ the code distance $d=9$.
-The error bars indicate one standard deviation.
+Figure 6: Logical error rates under the depolarizing noise and $N_{\mathrm{SA}}=1,\dots,10$ for the code distance $d=9$. The error bars indicate one standard deviation.
+Figure 7: Logical error rates under $Y$-biased noise where $p_{x}:p_{y}:p_{z}=1:5:1$ and $N_{\mathrm{SA}}=1,\dots,100$ for the code distance $d=9$. The error bars indicate one standard deviation.
+Figure 8: Logical error rates under $Y$-biased noise where $p_{x}:p_{y}:p_{z}=1:10:1$ and $N_{\mathrm{SA}}=1,\dots,100$ for the code distance $d=9$. The error bars indicate one standard deviation.
 In Figs.
-68, we observe that the logical error rate $P_{\mathrm{L}}$ of our SA decoder monotonically decreases as $N_{\mathrm{SA}}$ increases.
+68, we observe that the logical error rate $P_{\mathrm{L}}$ of our SA decoder decreases monotonically as $N_{\mathrm{SA}}$ increases.
 For $N_{\mathrm{SA}}\geq 2$, our SA decoder achieves logical error rates lower than our greedy matching decoder and is therefore an improvement over our greedy matching.
-For $N_{\mathrm{SA}}\geq 3$, our SA decoder is more accurate than the MWPM decoder.
+For $N_{\mathrm{SA}}\geq 3$, our SA decoder achieves a lower logical error rate than the MWPM decoder.
 To achieve the accuracy by the CPLEX decoder, under the depolarizing noise in Fig.
@@
 8, $N_{\mathrm{SA}}=100$ is not sufficient.
-Hence, to attain the optimal logical error rate, we observe that the more strongly $Y$-biased noise, the more $N_{\mathrm{SA}}$ our SA decoders need.
-V.3 Convergence of the SA decoders
-To construct a pure error, there are the following three methods:
+Hence, to attain the optimal logical error rate, we observe that the more strongly $Y$-biased the noise is, the larger $N_{\mathrm{SA}}$ our SA decoder needs.
+V.3 Effect of initial error configuration on SA performance
+To construct a reference configuration $R(S)$, we consider the following three methods:
 Boundary
-Every incorrect syndrome is directly connected with closer one in the top and bottom boundaries for white faces (the left and right boundaries for gray faces).
-We employ the SA decoder described in Algorithm 1.
+Each syndrome defect is connected to the nearer of the two boundaries (top or bottom for white faces; left or right for gray faces).
+We call this reference configuration construction routine, instead of the greedy_matching function in Algorithm 1.
 Greedy matching (same)
-It is Algorithm 5 invoked in the SA decoder of Algorithm 1, described in Sec. IV.3. For every invocation, it produces the same pure error.
+We use Algorithm 5 as the initializer in Algorithm 1; it returns the same reference configuration on every call.
 Greedy matching (different)
@@
 IV.4.
-For every invocation, it may produce a different pure error.
-We use this method in the present paper, unless otherwise stated.
+For every invocation, it may produce a different reference configuration.
+We use this method in this paper, unless otherwise stated.
 To compare the effect of these methods on the convergence of our SA decoder, we evaluate the ratio of logical error rates
@@
 where $P^{\mathrm{(SA)}}_{\mathrm{L}}$ and $P^{\mathrm{(CPLEX)}}_{\mathrm{L}}$ are the logical error rates obtained by the SA and CPLEX decoders, respectively.
-Here, the power $2/(d+1)$ is to cancel out $d$-dependence.
-As for the ratio $R_{\mathrm{L}}$, the lower value implies that our SA decoder is more accurate than the CPLEX decoder, and the value $1$ indicates that our SA decoder attains CPLEX decoder’s accuracy.
+Here, the exponent $2/(d+1)$ is introduced to cancel out $d$-dependence.
+A value $R_{\mathrm{L}}=1$ indicates that the SA decoder matches the CPLEX benchmark.
+Values above $1$ indicate worse performance than CPLEX.
 Fig.
@@
 In Figs.
-911, the ratio values $R_{\mathrm{L}}$ obtained by the greedy matching and MWPM decoders are drawn by dotted gray and dashed-dotted brown horizontal lines, respectively.
+911, the ratio values $R_{\mathrm{L}}$ obtained by the greedy matching and MWPM decoders are drawn by dotted gray and dash-dotted brown horizontal lines, respectively.
 (a) Refer to caption(b) Refer to caption (c) Refer to caption
-Figure 9: Convergence of our SA decoder for different construction methods of pure errors under the depolarizing noise.
+Figure 9: Convergence of our SA decoder for different construction methods of reference configurations under the depolarizing noise.
 The horizontal axis is the number of inverse temperatures $N_{\beta}$, and the vertical axis is the ratio $R_{\mathrm{L}}$.
@@
 (a) $p=0.04$ (b) $p=0.1$ (c) $p=0.15$.
+Dotted gray and dash-dotted brown horizontal lines show the ratios obtained by the greedy matching and MWPM decoders, respectively.
 (a) Refer to caption(b) Refer to caption (c) Refer to caption
-Figure 10: Convergence of our SA decoder for different construction methods of pure errors under the $Y$-biased noise such that $p_{x}:p_{y}:p_{z}=1:5:1$.
+Figure 10: Convergence of our SA decoder for different construction methods of reference configurations under the $Y$-biased noise such that $p_{x}:p_{y}:p_{z}=1:5:1$.
 The horizontal axis is the number of inverse temperatures $N_{\beta}$, and the vertical axis is the ratio $R_{\mathrm{L}}$.
@@
 (a) $p=0.04$ (b) $p=0.1$ (c) $p=0.15$.
+Dotted gray and dash-dotted brown horizontal lines show the ratios obtained by the greedy matching and MWPM decoders, respectively.
 (a) Refer to caption(b) Refer to caption (c) Refer to caption
-Figure 11: Convergence of our SA decoder for different construction methods of pure errors under the $Y$-biased noise such that $p_{x}:p_{y}:p_{z}=1:10:1$.
+Figure 11: Convergence of our SA decoder for different construction methods of reference configurations under the $Y$-biased noise such that $p_{x}:p_{y}:p_{z}=1:10:1$.
 The horizontal axis is the number of inverse temperatures $N_{\beta}$, and the vertical axis is the ratio $R_{\mathrm{L}}$.
@@
 (a) $p=0.04$ (b) $p=0.1$ (c) $p=0.15$.
+Dotted gray and dash-dotted brown horizontal lines show the ratios obtained by the greedy matching and MWPM decoders, respectively.
 In Figs.
-911, under all the noises and for all the pure error construction methods, we can see that the ratio $R_{\mathrm{L}}$ broadly decreases to the value $1$ as $N_{\beta}$ increases.
+911, under all the noises and for all the reference configuration construction methods, we see that the ratio $R_{\mathrm{L}}$ broadly decreases to the value $1$ as $N_{\beta}$ increases.
 For $N_{\beta}\geq 20$, the ‘Greedy matching (same)’ and ‘Greedy matching (different)’ methods outperform the MWPM decoder.
-Setting $N_{\beta}=0$ disables SA.
-Hence, $N_{\mathrm{SA}}$ indicates the number of not SAs but pure errors.
-In this setting, the ‘Greedy matching (same)’ and ‘Greedy matching (different)’ decoders become their MLD variants.
+For the ‘Greedy matching (same)’ and ‘Greedy matching (different)’, when $N_{\beta}=0$, the simulated_annealing function performs no Metropolis updates, so the decoder compares only the energies of the initial configurations.
+We refer to these variants as “initial-configuration-only” variants.
 In all the cases (a), (b), (c) of Figs.
-911, the ‘Greedy matching (different)’ method with $N_{\beta}=0$ achieves higher accuracy than the MWPM decoder.
-For all values of $N_{\beta}$ in Figs.
-911, ‘Greedy matching (different)’ is the highest accurate, followed by ‘Greedy matching (same)’, while the ‘Boundary’ method is the least accurate.
-The causes of these accuracy differences are considered as follows:
-  • Whereas the ‘Boundary’ method ignores the actual error probabilities $p_{x}$, $p_{y}$, $p_{z}$, the ‘Greedy matching (same)’ and ‘Greedy matching (different)’ methods take $p_{x}$, $p_{z}$ (but not $p_{y}$) into account.
-Consequently, the initial error configuration produced by ‘Greedy matching (same)’ is typically more similar to the actual error configuration than that produced by the ‘Boundary’ method.
-Thus, the ‘Greedy matching (same)’ method is more accurate than the ‘Boundary’ method.
-  • The ‘Greedy matching (different)’ method may produce a different initial error configuration for each SA.
-It leads to numerous variations of initial configurations, increasing the chance of reaching lower-energy states and thereby reducing the logical error rate.
+911, the initial-configuration-only variant of the ‘Greedy matching (different)’ method achieves higher accuracy than the MWPM decoder.
+In Figs.
+911, for all plotted values of $N_{\beta}$ and for all noise biases, ‘Greedy matching (different)’ consistently yields the lowest $R_{\mathrm{L}}$ among the three methods, followed by ‘Greedy matching (same)’, while the ‘Boundary’ method performs worst.
+We attribute these accuracy differences to the following factors:
+  • The ‘Boundary’ method ignores the error probabilities $p_{x}$, $p_{y}$, and $p_{z}$, whereas ‘Greedy matching (same)’ incorporates $p_{x}$ and $p_{z}$ through the weights $\tilde{\alpha}_{x}$ and $\tilde{\alpha}_{z}$.
+As a result, ‘Greedy matching (same)’ yields a bias-informed starting configuration.
+  • The ‘Greedy matching (different)’ method randomizes the tie-breaking among equal-distance pairs, generating diverse initial configurations across runs.
+This diversity increases the chance of reaching lower-energy configurations and thereby reduces the logical error rate.
 V.4 Comparison of decoding time with other decoding algorithms
@@
 From Figs.
-1214, we can see that the larger $d$, the longer time both the SA and CPLEX decoders take.
+1214, we see that the larger $d$ is, the longer both the SA and CPLEX decoders take.
 (a) Refer to caption (b) Refer to caption (c) Refer to caption
 Figure 12: Elapsed times for the SA and CPLEX decoders under the depolarizing noise.
-The vertical axis is logarithmic scale.
+The vertical axis denotes elapsed time with logarithmic scale.
+The orange dashed line shows the elapsed time of SA divided by $4N_{\mathrm{SA}}$, an idealized runtime estimate assuming perfect parallel efficiency across $4N_{\mathrm{SA}}$ independent SA instances and negligible overhead.
 (a) $p=0.02$ (b) $p=0.1$ (c) $p=0.15$.
-In (a) and (b), the dashed orange line implies elapsed time divided by $4N_{\mathrm{SA}}$, which is time assuming parallel execution of every SA.
-In (c), the CPLEX decoder indicates exponential time for the code distance $d$.
+In (c), the CPLEX decoder exhibits exponential scaling for the code distance $d$.
 (a) Refer to caption (b) Refer to caption (c) Refer to caption
 Figure 13: Elapsed times for the SA and CPLEX decoders under the $Y$-biased noise where $p_{x}:p_{y}:p_{z}=1:5:1$.
-The vertical axis is logarithmic scale.
+The vertical axis is on a logarithmic scale.
+The orange dashed line shows the elapsed time of SA divided by $4N_{\mathrm{SA}}$, an idealized runtime estimate assuming perfect parallel efficiency across $4N_{\mathrm{SA}}$ independent SA instances and negligible overhead.
 (a) $p=0.02$ (b) $p=0.1$ (c) $p=0.15$.
-In (a) and (b), the dashed orange line implies elapsed time divided by $4N_{\mathrm{SA}}$, which is time assuming parallel execution of every SA.
-In (c), the CPLEX decoder indicates exponential time for the code distance $d$.
+In (c), the CPLEX decoder exhibits exponential scaling for the code distance $d$.
 (a) Refer to caption (b) Refer to caption (c) Refer to caption
 Figure 14: Elapsed times for the SA and CPLEX decoders under the $Y$-biased noise where $p_{x}:p_{y}:p_{z}=1:10:1$.
-The vertical axis is logarithmic scale.
+The vertical axis is on a logarithmic scale.
+The orange dashed line shows the elapsed time of SA divided by $4N_{\mathrm{SA}}$, an idealized runtime estimate assuming perfect parallel efficiency across $4N_{\mathrm{SA}}$ independent SA instances and negligible overhead.
 (a) $p=0.02$ (b) $p=0.1$ (c) $p=0.15$.
-In (a) and (b), the dashed orange line implies elapsed time divided by $4N_{\mathrm{SA}}$, which is anticipated time assuming parallel execution of every SA.
-In (c), the CPLEX decoder indicates exponential time for the code distance $d$.
+In (c), the CPLEX decoder exhibits exponential scaling for the code distance $d$.
 For each decoder, we examine the dependence on $d$ and $p$.
-For all the physical error rates $p$, our SA decoder requires the same time, because the computational cost is determined solely by $N_{\beta}$ and $d$.
+For all the physical error rates $p$, our SA decoder takes approximately the same time, because the iteration counts are determined solely by $N_{\beta}$ and $d$ and does not depend on the syndrome weight $|S|$.
 Recalling that for each inverse temperature, we perform the Metropolis step $\mathcal{O}(N_{\mathrm{stabilizer}})=\mathcal{O}(d^{2})$ times, the total elapsed time scales accordingly.
 On the other hand, the CPLEX decoder exhibits exponential scaling in computation time with respect to the code distance $d$.
-This exponentiality is clearly visible for the case (c) $p=0.15$ in Figs.
+This exponential behavior is clearly visible for the case (c) $p=0.15$ in Figs.
 1214, where it is prohibitive to calculate for $d>11$.
-Let us compare the elapsed times of our SA decoder and the CPLEX decoder.
+Let us now compare the elapsed times of our SA decoder and the CPLEX decoder.
 For $p=0.15$, our SA decoder is clearly faster than the CPLEX decoder, whereas for $p=0.02,0.1$, it requires more time than the CPLEX decoder.
 In each of Figs.
-1214, ‘SA (divided by $4N_{\mathrm{SA}}$)’ drawn with the dashed orange line is the elapsed time divided by $4N_{\mathrm{SA}}=40$, indicating anticipated time in parallelized execution.
-By comparing the dashed orange line and the blue line, we can see that our SA decoder is faster than the CPLEX decoder.
-Consequently, if our SA decoder is parallelized, it outperforms the CPLEX decoder for all the values $d$ and $p$ we employ.
+1214, ‘SA (divided by $4N_{\mathrm{SA}}$)’ drawn with the orange dashed line shows the elapsed time divided by $4N_{\mathrm{SA}}=40$, which corresponds to an idealized runtime assuming perfect parallel efficiency (i.e., negligible overhead and ideal strong scaling).
+Comparing the idealized parallel estimate (orange dashed) with the CPLEX time (blue) suggests that, under near-perfect scaling across the $4N_{\mathrm{SA}}$ independent runs, SA could be faster than CPLEX over the tested $(d,p)$ range.
+Quantifying the achievable parallel efficiency and overhead is left for future work.
 Next, we present the elapsed times for various decoding algorithms in Fig. 15.
 […]
-For our SA decoder, we set the number of inverse temperatures $N_{\mathrm{\beta}}=100$ and $N_{\mathrm{SA}}=10$ in Algorithm 7.
-Note that $N_{\mathrm{SA}}$ can be any number, because every SAs have almost the same computational cost.
-For the MPS decoder, we used the implementation for the CSS surface code in the repository [39], because the decoder for the XZZX code employed in Ref.
-[9] is not open to the public.
+For our SA decoder, we set the number of inverse temperatures $N_{\beta}=100$ and $N_{\mathrm{SA}}=10$ in Algorithm 7.
+Note that increasing $N_{\mathrm{SA}}$ increases the total cost roughly linearly, since each SA run has a similar per-run cost.
+The MPS timing is provided only as a rough reference: we use the CSS surface-code implementation in Ref.
+[qecsim], since the XZZX-specific implementation used in Ref.
+[xzzx_nature] is not publicly available.
 Here, we employ the truncation dimension $\chi=8$ for the MPS decoder.
@@
 15, we observe that our SA decoder is faster than the MPS decoder except for $d=9$.
-This is because our SA decoder is implemented in C++, whereas the MPS decoder is implemented in Python.
-As mentioned in Ref.
-[17, 8], the MPS decoder has a computational complexity of $\mathcal{O}(N_{q})=\mathcal{O}(d^{2})$, which is the same as our SA decoder.
+This is partially because our SA decoder is implemented in C++, whereas the MPS decoder is implemented in Python.
+For a fixed truncation dimension $\chi$, the MPS decoder scales as $\mathcal{O}(N_{\mathrm{q}})=\mathcal{O}(d^{2})$ (up to polynomial factors in $\chi$[efficient_MLD, tuckett_tailoring].
+For fixed $N_{\beta}$ and $N_{\mathrm{SA}}$, our SA decoder also scales as $\mathcal{O}(d^{2})$ in $d$, although constant factors and parameter dependencies differ.
 Our SA decoder labeled by ‘SA’ is slower than the CPLEX decoder.
 However, it can be readily parallelized.
-Thus, we can consider the elapsed time divided by $4N_{\mathrm{SA}}$, denoted by ‘SA (divided by $4N_{\mathrm{SA}}$)’, again, which takes less time than ‘CPLEX’.
-Consequently, among the decoders which can deal with the error probability $p_{y}$, such as the SA, CPLEX and MPS decoders, our SA decoder is the fastest if parallelized.
+Thus, dividing the ‘SA’ time by $4N_{\mathrm{SA}}$ yields the idealized parallel runtime, denoted by ‘SA (divided by $4N_{\mathrm{SA}}$)’, which is lower than that of ‘CPLEX’.
+Consequently, among the decoders which can deal with the error probability $p_{y}$, such as the SA, CPLEX and MPS decoders, our SA decoder would be the fastest under an idealized perfect parallelization assumption.
 From Fig.
@@
 IV.3, which is lower than the worst-case complexity of the MWPM decoder $\mathcal{O}(|S_{\bullet}|^{3}\log{|S_{\bullet}|})=\mathcal{O}(d^{6}\log{d})$ in Ref.
-[37].
-Note that whereas in circuit-level noise model, the MWPM decoder uses the Dijkstra algorithm to find a path with minimum weight between any possible pair of incorrect syndromes, in the code capacity noise model, it does not because the weights are explicitly given by Eqs.
-(23),(24) or (25).
-Note that the MWPM decoder has an average computational complexity of $\mathcal{O}(d^{2})$ from numerical simulation [35].
-Investigating the average complexity for our greedy matching decoder would be meaningful, and is left to our future work.
+[BlossomV]  [^6] Note that, in circuit-level noise model, the MWPM decoder is used with the Dijkstra algorithm to find a path with a minimum weight between any possible pair of syndrome defects, whereas in the code capacity noise model, the Dijkstra algorithm is not needed, because the weights are explicitly given by Eqs.
+(26), (28), (29), or (30)..
+We emphasize that practical MWPM decoders often show much better average-case scaling (e.g., $\mathcal{O}(d^{2})$ reported in Ref.
+[MWPM_complexity]), and a similar empirical study for our greedy matching decoder is left for future work.
 VI Discussion
@@
 V.3, under much more strongly $Y$-biased noise, our SA decoder requires a large $N_{\mathrm{SA}}$ in order to attain the optimal logical error rate.
-A promising remedy to this problem is to modify the definition of stabilizer generators to include $Y$ operators, analogously to the tailored code for the CSS surface code [8].
-Specifically, for the case $p_{y}\geq p_{z}\geq p_{x}$, we should employ ‘XYYX code’.
-This modified code can detect the endpoints of error chains which consist of $Z$ ($Y$) errors on horizontal (vertical) edges, respectively.
+A possible remedy is to modify the definition of stabilizer generators, analogously to the tailored code for the CSS surface code [tuckett_tailoring].
+For example, in regimes where $p_{y}$ dominates (e.g., $p_{y}\geq p_{z}\geq p_{x}$), one could consider an “XYYX” variant whose stabilizer generators include $Y$ operators.
+This modified code can detect the endpoints of error chains which consist of $Z$ and $Y$ errors on horizontal and vertical edges, respectively.
 Thus, in our greedy matching algorithm, we can take into account the probabilities $p_{z}$ and $p_{y}$.
-We therefore expect to obtain a pure error, closely resemblant to the actual recovery configuration.
-By employing this ‘good’ initial configuration, our SA decoder is expected to quickly converge to the optimal logical error rate.
+We therefore expect to obtain a reference configuration, closely resembling the true error configuration.
+By employing this better initial configuration, our SA decoder is expected to converge more quickly to the optimal logical error rate.
 Such a treatment can be applied to accommodate the given noise bias.
 A key advantage of our SA decoder is that it provides explicit control over the trade-off between computational cost and accuracy by adjusting $N_{\beta}$.
-That is, increasing the number of inverse temperatures $N_{\beta}$ improves the decoding accuracy, but also increases the computational cost.
+That is, increasing the number of inverse temperatures $N_{\beta}$ typically improves the decoding accuracy, but also increases the computational cost.
 In contrast, the cost of deterministic algorithms such as the MWPM, CPLEX, and MPS decoders cannot be easily adjusted, and their parallelization is neither straightforward nor efficient.
 VII Conclusion
-We have demonstrated our decoder for the XZZX code by using simulated annealing (SA).
-As an initial configuration of SA, we have proposed to employ recovery chains obtained by our decoder which utilizes a kind of greedy matching.
-We assumed the code capacity noise, where only the data qubits suffer errors. Under the depolarizing noise and the $Y$-biased noises, we performed the numerical simulation.
-We observed that our SA decoder achieves the logical error rate $P_{\mathrm{L}}$, equal to the optimal rate by the CPLEX decoder, which is formulated as an integer programming approach.
+We have formulated and demonstrated our SA decoder for the XZZX code.
+As an initial configuration for SA, we have proposed to employ recovery chains obtained by our decoder that utilizes a greedy matching graph algorithm.
+We assumed the code capacity noise model. Under this model, we performed numerical simulations under depolarizing noise and $Y$-biased Pauli noise.
+We observed that our SA decoder achieves the logical error rates $P_{\mathrm{L}}$, comparable to those of the CPLEX decoder, which we use as an optimal MAP benchmark.
 Moreover, it is more accurate than the MWPM decoder, which cannot take into account the probability of $Y$ errors.
-We examined the convergence of the logical error rate for the number of inverse temperatures $N_{\mathrm{\beta}}$ we employ for our SA decoder.
-In particular, our SA decoder can be interpreted as giving improvement of initial configuration by our greedy matching decoder.
-Furthermore, we confirmed that the convergence is strongly influenced by methods of constructing a pure error, which are employed to prepare initial configurations.
-Especially, we proposed the ‘Greedy matching (different)’ method which may produce a different pure error for every invocation, and observed that it leads to markedly faster convergence.
+We examined the convergence of the logical error rate for the number of inverse temperatures $N_{\beta}$ we employ for our SA decoder.
+In particular, our SA decoder improves upon the initial configuration obtained from our greedy matching decoder.
+Furthermore, we confirmed that the convergence is strongly influenced by methods of constructing a reference configuration, which are employed to prepare initial configurations.
+Especially, we proposed the ‘Greedy matching (different)’ method which may produce a different reference configuration for every invocation, and observed that it leads to markedly faster convergence.
 Subsequently, we compared the elapsed times of various decoding algorithms.
-As a result, our SA decoder is faster than the CPLEX decoder when the physical error rate $p$ is higher.
-For lower $p$, our SA decoder, if parallelized, is faster than the CPLEX decoder.
-In summary, we conclude that our parallelizable SA decoder is the fastest among the decoders, including the MPS decoder, which can take into account the probability of $Y$ errors.
-This parallelization of SAs and its simulation remains a topic for our future work.
-We list other future works below. Although our analysis has assumed the code capacity noise model, applying our SA decoder to phenomenological noise and circuit-level noise models is straightforward.
+As a result, our SA decoder is faster than the CPLEX decoder when the physical error rate $p$ is relatively high.
+In the lower-$p$ regime, our SA decoder, under an idealized perfect parallelization assumption, is faster than the CPLEX decoder.
+In summary, we conclude that our parallelizable SA decoder would be the fastest among the decoders that can take into account the probability of $Y$ errors.
+Quantifying practical parallel efficiency and overhead remains for future work.
+We list other topics for future work below.
+Although our analysis has assumed the code capacity noise model, extending our SA decoder to phenomenological noise and circuit-level noise models is conceptually straightforward.
 The decoding time may be reduced by improving the MCMC algorithm.
-Our SA decoder consists of independent SA processes, making it well-suited for parallelized implementation in FPGA (field-programmable gate array).
-Alternatively, techniques to accelerate the MCMC method could be employed, such as the replica-exchange Monte Carlo method [40] and cluster algorithms including the Wolff algorithm [41].
-Furthermore, the Metropolis algorithm can be parallelized by domain (area) decomposition, which is implementable in FPGA [42].
-In our SA decoder, we are able to control the trade-off between the decoding accuracy and time, by adjusting the number of Monte Carlo steps, temperature scheduling and convergence criterion, depending on the code distance $d$ and the physical error rate $p$.
-Finding a suitable set of parameter values may require exhaustive investigation.
-The MCMC decoders require only the definitions of stabilizer generators and logical operators of codes, without relying on a decoding graph and, hence, they are applicable to many types of codes, including the color code and quantum LDPC codes [43].
-To obtain a pure error to construct an initial error configuration, we can use not only our greedy matching decoder but also other decoders, such as the union-find decoder [44].
-For faster convergence, we will pursue ‘good’ initial error configuration.
-We plan to perform simulations under actual noise biases $p_{x}:p_{y}:p_{z}$ specific to individual quantum devices.
-Toward practical use in quantum computing, it is necessary to improve the decoding speed, particularly at lower physical error rates $p$.
-Our results suggest that our SA decoder is a promising candidate for practical quantum computing.
+Our SA decoder consists of independent SA processes, making it well-suited for parallelized implementation on FPGA (field-programmable gate array).
+Alternatively, techniques to accelerate the MCMC method could be employed, such as the replica-exchange Monte Carlo method [hukushima] and cluster algorithms including the Wolff algorithm [Wolff_algorithm].
+Furthermore, the Metropolis algorithm can be parallelized by domain (area) decomposition, which is implementable on FPGA [parallel_FPGA].
+Our SA decoder allows an explicit trade-off between decoding accuracy and runtime through tunable parameters such as the number of sweeps, the temperature schedule, and the stopping criterion.
+A systematic study of parameter tuning as a function of $(d,p)$ will be worthwhile.
+An advantage of MCMC-based decoders is their generality: they require only the stabilizer generators and a specification of logical operators, without an explicit decoding graph.
+This makes them broadly applicable, e.g., to the color code and to quantum LDPC codes [quantum_LDPC_codes].
+For SA initialization, one may use reference configurations produced not only by our greedy matching decoder but also by other fast decoders, such as union-find [union-find_decoder].
+Several decoders have been proposed that explicitly account for correlations induced by $Y$ errors [fusion_blossom, fowler2013optimalcomplexitycorrectioncorrelated, Paler_2023].
+A systematic comparison with these approaches will be an important direction for future work.
+Finally, we plan to perform simulations under device-specific Pauli-noise biases $p_{x}:p_{y}:p_{z}$.
+Toward practical deployment, further speed improvements will be particularly important in the low-$p$ regime.
 Acknowledgements.
-We would like to thank Keisuke Fujii for his helpful discussions.
+We would like to thank Yusaku Takeuchi, Yugo Takada, Mitsuki Katsuda, Hiroshi Ueda, Hideaki Hakoshima, Kosuke Mitarai, Jun Fujisaki, Hirotaka Ohshima, Shintaro Sato, and Keisuke Fujii for their helpful discussions.
 This work is supported by MEXT Quantum Leap Flagship Program (MEXT Q-LEAP) Grant No.
-JPMXS0118067394 and JPMXS0120319794, JST COI-NEXT Grant No.
-JPMJPF2014, and JST Moonshot R&D Grant No.
-JPMJMS2061.
+JPMXS0118067394 and JPMXS0120319794; JST Moonshot R&D Grant No.
+JPMJMS2061; JST COI-NEXT projects, Quantum Software Research Hub (Grant No.
+JPMJPF2014) and the Center of Innovation for Sustainable Quantum AI (Grant No.
+JPMJPF2221).
 Appendix A Decoder formulated as integer programming problem
-In this section, instead of Eq. (20), we formulate decoding problem as finding the recovery configuration $C$, subject to the constraint $\partial{C}=S$:
+In this section, instead of Eq. (21), we formulate the decoding problem as finding the recovery configuration $C$, subject to the constraint $\partial{C}=S$:
 $\arg\min_{C}\left\{\mathcal{H}(C)\,\middle|\,\partial{C}=S\right\}.$ (38)
-The error configuration $C$ can be represented as the set of one bit binary variables $\{x_{i},y_{i},z_{i}\}_{i}$ where $i=1,\dots,N_{\mathrm{q}}$ indexes the qubits.
-Namely, $x_{i}=1$ indicates that the qubit $i$ suffers $X$ error, while the variables $y_{i}$ and $z_{i}$ similarly represent $Y$ and $Z$ errors, respectively.
-Using this binary representation, the syndrome consistency condition $\partial{C}=S$ in Eq.
-(33) is explicitly expressed as a system of equations: For each face $f$,
+The error configuration $C$ can be represented as the set of one-bit binary variables $\{x_{i},y_{i},z_{i}\}_{i}$ with the condition $x_{i}+y_{i}+z_{i}\leq 1$ where $i=1,\dots,N_{\mathrm{q}}$ indexes the qubits.
+Namely, $x_{i}=1$ indicates that the qubit $i$ suffers an $X$ error, while the variables $y_{i}$ and $z_{i}$ similarly represent a $Y$ error and a $Z$ error, respectively.
+Using this binary representation, the syndrome-consistency condition $\partial{C}=S$ in Eq.
+(38) is explicitly expressed as a system of equations.
+That is, for each face $f$,
 $\begin{split}\left(z_{\mathrm{left}(f)}\oplus y_{\mathrm{left}(f)}\right)\oplus\left(z_{\mathrm{right}(f)}\oplus y_{\mathrm{right}(f)}\right)&\\ \,\oplus\left(x_{\mathrm{up}(f)}\oplus y_{\mathrm{up}(f)}\right)\oplus\left(x_{\mathrm{down}(f)}\oplus y_{\mathrm{down}(f)}\right)&=s_{f}\end{split}$ (39)
-where $s_{f}\in\{0,1\}$ is the syndrome measurement outcome, as defined in Eq.
-(2), and the operation $\oplus$ denotes addition modulo two.
+where $s_{f}\in\{0,1\}$ is the syndrome measurement outcome determined in Eq.
+(2), and the operation $\oplus$ denotes XOR (addition modulo two) [^7] To solve Eq.
+(38) with CPLEX, we convert the parity (XOR) constraints in Eq.
+(39) into an equivalent integer linear programming using auxiliary integer variables and linear constraints.
+.
+For boundary faces, where the stabilizer has weight three, the missing-qubit term is omitted from the parity (XOR) constraint.
 Under the constraint Eq.
@@
 This solution yields a recovery configuration $C$.
-This problem falls into the category of (mixed-) integer linear programming, known to be NP-hard.
+This problem falls into the category of binary integer programming, known to be NP-hard.
 The integer programming formulation exhibits exponential computational complexity with respect to both the code distance $d$ and the physical error rate $p$.
-Thus, for large values of $d$ or $p$, it is prohibitive to find a solution in a realistic time.
-To solve the integer linear programming problem, we employ the C++ library, CPLEX (version 22.1.0) [45], whence we call it CPLEX decoder.
-The use of integer programming solvers for decoding was originally proposed and demonstrated for the color code [46].
-Since the CPLEX decoder can take the probabilities $p_{x}$, $p_{y}$, $p_{z}$ into account by Eq.
-(10), it is capable of achieving the lowest logical error rate among the decoders formulated as the energy minimization problems.
-Thus, we regard the logical error rate obtained by the CPLEX decoder as a reliable benchmark and thereby evaluate the accuracy of our SA decoder in Sec.
-V.
+Thus, for large values of $d$ or $p$, it is prohibitive to find a solution within a reasonable amount of time.
+To solve the integer linear programming problem, we employ the C++ library, CPLEX (version 22.1.0) [cplex], which we call the CPLEX decoder.
+The use of integer programming solvers for decoding was originally proposed and demonstrated for the color code [color_code_integer].
+We use the CPLEX decoder as an exact benchmark for the minimum-energy (MAP-configuration) objective $\mathcal{H}(C)$, which includes the probabilities $p_{x}$, $p_{y}$, $p_{z}$.